### v2

# An average Brun–Titchmarsh theorem and shifted primes with a large prime factor

Abstract.
The author studies an average version of Brun–Titchmarsh theorem with large moduli. Using Maynard’s recent breakthrough on the Bombieri–Friedlander–Iwaniec type triple convolution estimates, we refine the previous result of Baker and Harman (1996). As an application, we improve a result of Baker and Harman (1998) on shifted primes with a large prime factor, showing that the largest prime factor of p1p-1 is larger than $p^{0.679}$ for infinitely many primes $p$.
Key words and phrases:
Prime, Sieve methods, Large prime factors
2020 Mathematics Subject Classification:
11N05, 11N35, 11N36
1. Introduction
One of the most important problems in number theory is the twin prime conjecture, which states that there are infinitely many primes $p$ such that $p+2$ is also prime. This conjecture still remains open, and there are three main ways to attack this conjecture. The first way is to study $p+2$ with few primary factors, and the best result is due to Chen [8], who proved in 1973 that there are infinitely many primes $p$ such that $p+2$ has at most two prime factors. One can also see [7], [36], [25], [24], [30] and the author’s preprints [20] [21] for further refinements. The second way is to study bounded gaps between consecutive primes, and we refer the interested readers to the works of Zhang [37], Maynard [27], Polymath8b [31] and Stadlmann [34].
The third way to attack the twin prime conjecture is to consider the largest prime factor of $p+2$. Let $P^{+}(n)$ denote the largest prime factor of $n$. Since we have $P^{+}(p)=p$, the twin prime conjecture is equivalent to the following statement: There are infinitely many $p$ such that
$P^{+}(p+2)>P^{+}(p)=p.$ (1)
Or, in general, there are infinitely many $p$ such that
$P^{+}(p+a)>p$ (2)
for a fixed non-zero integer $a$. We don’t know how to prove (2), but we can prove some weaker results along this way. Using the Bombieri–Vinogradov Theorem proved in 1965, Goldfeld [14] showed in 1969 that there are infinitely many $p$ such that
$P^{+}(p+a)>p^{\boldsymbol{\Theta}-\varepsilon}$ (3)
for some $\boldsymbol{\Theta}>\frac{1}{2}$. After Goldfeld, many mathematicians improved the lower bound for $\boldsymbol{\Theta}$. Let $\theta\in\left[\frac{1}{2},\boldsymbol{\Theta}\right]$. In order to deal with $\theta$ larger than $\frac{1}{2}$, one needs to study an average version of Brun–Titchmarsh inequality, and prove strong upper estimates
$\pi\left(x;q,a\right)\leqslant\frac{C(\theta)x}{\varphi(q)\log x}$ (4)
with an upper constant $C(\theta)$ for almost all moduli $q\leqslant x^{\theta}$ (with at most $O\left(x^{\theta}(\log x)^{-A}\right)$ exceptions). The progression of records for $\boldsymbol{\Theta}$ can be seen in the following table.
$\boldsymbol{\Theta}$ $\approx$ Author Year New ”almost all” Brun–Titchmarsh result
$1-\frac{1}{2}e^{-\frac{1}{4}}$ $0.6105$ Motohashi [29] 1970 $\pi\left(x;q,a\right)\leqslant\frac{\left(\frac{2}{1-\theta}+\varepsilon\right)x}{\varphi(q)\log x},\ q<x,\ q\sim x^{\theta}\ (\text{for all }q)$
$2-\frac{3}{2}e^{-\frac{1}{12}}$ $0.6199$ Hooley [17] 1972
[[17], Theorem 1]
$\pi\left(x;q,a\right)\leqslant\frac{\left(\frac{6}{2-\theta}+\varepsilon\right)x}{\varphi(q)\log x},\ x^{\frac{1}{2}}<q<x^{\frac{4}{5}},\ q\sim x^{\theta}$
$\frac{5}{8}$ $0.625$ Hooley [18] 1973
[[9], (2)]
$\pi\left(x;q,a\right)\leqslant\frac{(4+\varepsilon)x}{\varphi(q)\log x},\ x^{\frac{1}{2}}<q<x^{\frac{3}{4}}$
$\frac{5}{3}-\frac{17^{\frac{5}{7}}}{2^{\frac{4}{7}}3^{\frac{5}{7}}5^{\frac{3}{7}}}e^{-\frac{1}{8}}$ $0.6381$ Iwaniec [19] 1982
[[19], Theorems 6 and 10]
$\pi\left(x;q,a\right)\leqslant\begin{cases}\frac{\left(\frac{8}{6-7\theta}+\varepsilon\right)x}{\varphi(q)\log x},&x^{\frac{1}{2}}<q<x^{\frac{8}{15}},\ q\sim x^{\theta}\ (\text{for all }q)\\ \frac{\left(\frac{12}{5-3\theta}+\varepsilon\right)x}{\varphi(q)\log x},&x^{\frac{8}{15}}<q<x^{\frac{2}{3}},\ q\sim x^{\theta}\end{cases}$
$1-\frac{1}{2}e^{-\frac{3}{8}}$ $0.6563$ Deshouillers and Iwaniec [9] 1984
[[9], Theorem]
$\pi\left(x;q,a\right)\leqslant\frac{\left(\frac{4}{3(1-\theta)}+\varepsilon\right)x}{\varphi(q)\log x},\ x^{\frac{1}{2}}<q<x,\ q\sim x^{\theta}$
$\frac{1}{20}\exp\left(\frac{-3}{280}\times\right$.
$\left(35-21\log\left(\frac{13}{12}\right)\right$.
$60\log\left(\frac{400}{351}\right)$
$\left.\left.+35\log\left(\frac{361}{360}\right)\right)\right)\times$
[…]
$\left(35-21\log\left(\frac{13}{12}\right)\right$.
$60\log\left(\frac{400}{351}\right)$
$\left.\left.\left.+35\log\left(\frac{361}{360}\right)\right)\right)-9\right)$
$0.6578$ Fouvry [11] 1984
[[11], Théorème]
$\pi\left(x;q,a\right)\leqslant\begin{cases}\frac{\left(\frac{12}{25-40\theta}-\log\left(\frac{10(1-\theta)}{9\theta}\right)+\varepsilon\right)x}{\varphi(q)\log x},&x^{\frac{1}{2}}<q<x^{\frac{53}{104}},\ q\sim x^{\theta}\\ \frac{\left(\frac{48}{47-56\theta}-\log\left(\frac{10(1-\theta)}{9\theta}\right)+\varepsilon\right)x}{\varphi(q)\log x},&x^{\frac{53}{104}}<q<x^{\frac{10}{19}},\ q\sim x^{\theta}\\ \frac{\left(\frac{48}{47-56\theta}+\varepsilon\right)x}{\varphi(q)\log x},&x^{\frac{10}{19}}<q<x^{\frac{11}{20}},\ q\sim x^{\theta}\end{cases}$
$0.6683$ $0.6683$ Fouvry [12] 1985
[[12], Théorème 3]
$\pi\left(x;q,a\right)\leqslant\begin{cases}\frac{(1+\text{Sieve Integrals}+\varepsilon)x}{\varphi(q)\log x},&x^{\frac{1}{2}}<q<x^{\frac{17}{32}},\ q\sim x^{\theta}\\ \frac{\left(\frac{22}{15-13\theta}-\log\left(\frac{4(1-\theta)}{3\theta}\right)+\varepsilon\right)x}{\varphi(q)\log x},&x^{\frac{17}{32}}<q<x^{\frac{8}{15}},\ q\sim x^{\theta}\\ \frac{\left(\frac{10}{9-10\theta}-\log\left(\frac{4(1-\theta)}{3\theta}\right)+\varepsilon\right)x}{\varphi(q)\log x},&x^{\frac{8}{15}}<q<x^{\frac{4}{7}},\ q\sim x^{\theta}\\ \frac{\left(\frac{10}{9-10\theta}+\varepsilon\right)x}{\varphi(q)\log x},&x^{\frac{4}{7}}<q<x^{\frac{3}{5}},\ q\sim x^{\theta}\\ \frac{\left(\frac{8}{3-\theta}+\varepsilon\right)x}{\varphi(q)\log x},&x^{\frac{3}{5}}<q<x^{\frac{5}{7}},\ q\sim x^{\theta}\end{cases}$
In 1996, Baker and Harman [1] presented a mixed approach to this problem. They combined the traditional average Brun–Titchmarsh approach (with a refined version of [[12], Théorème 3] which will be used in Section 6.10) with the alternative sieve developed by Harman [15], using the estimates of Bombieri–Friedlander–Iwaniec type (see [5] and [6], for example) as arithmetical information inputs. Their method is more powerful when $\theta$ is near to $\frac{1}{2}$. In this way they got $\boldsymbol{\Theta}=0.676$. Two years later, they [2] further developed a variant of the three-dimensional alternative sieve used in [1] and obtained a saving over one-dimensional sums, leading to a better result $\boldsymbol{\Theta}=0.677$. Although this result only improved the previous one by $0.001$, it was still a improvement since they used a continuous version of alternative sieve, and any saving on a single point became quite small after integrating. Because of the lack of new arithmetical information, even $0.678$ seems impossible without any new things.
Recently, Maynard [28] got a new breakthrough on the distribution of primes in arithmetical progressions to large moduli $q=q_{1}q_{2}\sim x^{\frac{11}{21}-\varepsilon}$. The main theorem in his paper is not applicable here, but we find that one of his lemmas (Proposition 8.3) can be used here as a new arithmetical information input. By inserting this into the sieving process in [1] and [2], together with a careful numerical calculation for sieve integrals, we get the following sharper result.
Theorem 1.
There are infinitely many primes $p$ such that
$P^{+}(p+a)>p^{0.679-\varepsilon}.$
In order to prove our Theorem 1, we need to give good upper bounds for $C(\theta)$ in (4). Clearly this yields an improvement over the main (numerical) theorem for $C_{2}(\theta)$ in [1]. All the numerical values of integrals in Section 6 are calculated using Mathematica 14.
Throughout this paper, we always suppose that $\varepsilon$ is a sufficiently small positive constant and $x$ is sufficiently large. Let $\theta\in[0.5,0.679]$ and $Q=x^{\theta}$. The letter $p$, with or without subscript, is reserved for prime numbers. We define the boolean function as
$\texttt{Boole}(\textbf{X})=\begin{cases}1&\text{if }\textbf{X}\text{ is true},\\ 0&\text{if }\textbf{X}\text{ is false}.\end{cases}$ (5)
2. An outline of the proof
Let
$\mathcal{A}^{q}=\{n:n\sim x,\ n\equiv a(\bmod q)\},\quad\mathcal{B}^{q}=\{n:n\sim x,\ (n,q)=1\},$
$\mathcal{C}_{d}=\{n:nd\in\mathcal{C}\},\quad P(z)=\prod_{p<z}p,\quad S(\mathcal{C},z)=\sum_{\begin{subarray}{c}n\in\mathcal{C}\\ (n,P(z))=1\end{subarray}}1.$
To prove our Theorem 1, it suffices to show that
$\sum_{x^{0.679}\leqslant p\leqslant x+a}\pi(x;p,-a)\log p>\varepsilon x.$ (6)
Note that we have
$\sum_{p\leqslant x}\log(p+a)=$ $\ \sum_{p\leqslant x}\sum_{n|(p+a)}\Lambda(n)$
$=$ $\ \sum_{p_{1}\leqslant x}\sum_{p_{2}|(p_{1}+a)}\log p_{2}+o(x)$
$=$ $\ \sum_{p\leqslant x+a}\pi(x;p,-a)\log p+o(x)$ (7)
and
$\sum_{p\leqslant x}\log(p+a)\sim x.$ (8)
By the Prime Number Theorem, we have
$\sum_{p\leqslant x+a}\pi(x;p,-a)\log p>(1-\varepsilon)x.$ (9)
From the Bombieri–Vinogradov theorem and the Brun–Titchmarsh inequality, we have
$\sum_{p\leqslant x^{\frac{1}{2}}}\pi(x;p,-a)\log p<\left(\frac{1}{2}+\varepsilon\right)x.$ (10)
Thus, in order to show (6), we only need to establish that
$\sum_{x^{\frac{1}{2}}<p\leqslant x^{0.679}}\pi(x;p,-a)\log p<\left(\frac{1}{2}-3\varepsilon\right)x.$ (11)
We remark that one can also see [22] and previous articles on that topic to get the idea of using a continuous version of Harman’s alternative sieve to give an upper bound for a special sum over primes in a fixed range. By the definition of the sieve function and Prime Number Theorem, we have
$\sum_{p\in\mathcal{A}^{q}}1=S\left(\mathcal{A}^{q},(2x)^{\frac{1}{2}}\right)\quad\text{and}\quad S\left(\mathcal{B}^{q},(2x)^{\frac{1}{2}}\right)=(1+o(1))\frac{x}{\log x}.$ (12)
Thus, we need an estimate for $S\left(\mathcal{A}^{q},(2x)^{\frac{1}{2}}\right)$ that holds for almost all $q\sim Q$. Our aim is to show that the sparser set $\mathcal{A}^{q}$ contains the expected proportion of primes compared to the larger set $\mathcal{B}^{q}$, which requires us to decompose $S\left(\mathcal{A}^{q},(2x)^{\frac{1}{2}}\right)$ and prove asymptotic formulas of the form
$S\left(\mathcal{A}^{q},z\right)=(1+o(1))\frac{1}{\varphi(q)}S\left(\mathcal{B}^{q},z\right)$ (13)
for some parts of it, and drop the other parts. The dropped parts must be negative, since we want an upper bound. If we have, for almost all $q\sim Q$,
$S\left(\mathcal{A}^{q},(2x)^{\frac{1}{2}}\right)<C(\theta)\frac{x}{\varphi(q)\log x},$ (14)
then trivially
$\pi(x;q,a)<C(\theta)\frac{x}{\varphi(q)\log x}.$ (15)
Note that $C(\theta)$ is monotonic increasing ([[1], Theorem 1(i)]). Combining (15) with the bound
$\int_{0.5}^{0.679}C(\theta)d\theta<\frac{1}{2},$ (16)
we can easily deduce (11) and prove Theorem 1. Now we only need to prove (16) by giving upper bounds for the function $C(\theta)$. In Section 6 we shall consider it with $\theta$ lies in different intervals. Note that each $\theta$ corresponds to an independent sieve decomposition.
In the process of giving asymptotic formulas like (13) in Harman’s sieve, one needs the following types of arithmetical information (with all variables are dyadic constrained, and all coefficients are divisor-bounded):
$\sum_{q\sim Q}\gamma_{q}\sum_{m}a_{m}\sum_{mn_{1}\cdots n_{j}\in\mathcal{A}^{q}}1=(1+o(1))\frac{1}{\varphi(q)}\sum_{q\sim Q}\gamma_{q}\sum_{m}a_{m}\sum_{mn_{1}\cdots n_{j}\in\mathcal{B}^{q}}1,$ (Type-Ij)
$\sum_{q\sim Q}\gamma_{q}\sum_{m_{1}\cdots m_{j}\in\mathcal{A}^{q}}a_{1,m_{1}}a_{2,m_{2}}\cdots a_{j,m_{j}}=(1+o(1))\frac{1}{\varphi(q)}\sum_{q\sim Q}\gamma_{q}\sum_{m_{1}\cdots m_{j}\in\mathcal{B}^{q}}a_{1,m_{1}}a_{2,m_{2}}\cdots a_{j,m_{j}},$ (Type-IIj)
$\sum_{q\sim Q}\gamma_{q}\sum_{m,n}a_{m}b_{n}\sum_{mnl\in\mathcal{A}^{q}}1=(1+o(1))\frac{1}{\varphi(q)}\sum_{q\sim Q}\gamma_{q}\sum_{m,n}a_{m}b_{n}\sum_{mnl\in\mathcal{B}^{q}}1.$ (Type-I/II)
We can combine Type-I1 and Type-II information together to prove asymptotic formulas for sums of $S\left(\mathcal{A}_{p_{1}\cdots p_{n}},x^{\kappa_{0}}\right)$, and sometimes we can use only Type-II information to give asymptotic formulas for sums of $S\left(\mathcal{A}_{p_{1}\cdots p_{n}},p_{n}\right)$. For the two-dimensional or even three-dimensional alternative sieve, Type-I2 and Type-I3 information is required.
3. Type-II arithmetical information
In this section we give several results of the form
$\sum_{q\sim Q}\gamma_{q}\sum_{\begin{subarray}{c}m_{1}\cdots m_{j}\in\mathcal{A}^{q}\\ m_{i}\sim M_{i}\\ 1\leqslant i\leqslant j\end{subarray}}a_{1,m_{1}}a_{2,m_{2}}\cdots a_{j,m_{j}}=(1+o(1))\frac{1}{\varphi(q)}\sum_{q\sim Q}\gamma_{q}\sum_{\begin{subarray}{c}m_{1}\cdots m_{j}\in\mathcal{B}^{q}\\ m_{i}\sim M_{i}\\ 1\leqslant i\leqslant j\end{subarray}}a_{1,m_{1}}a_{2,m_{2}}\cdots a_{j,m_{j}}.$
The three conditions we want the coefficients to satisfy are as follows:
(Condition 3.1: Siegel–Walfisz condition) For any $f\geqslant 1$, $k\geqslant 1$, $b\neq 0$ and $(k,b)=1$, we have
$\sum_{\begin{subarray}{c}l\sim L\\ l\equiv b(\operatorname{mod}k)\\ (l,f)=1\end{subarray}}\lambda_{l}=\frac{1}{\varphi(k)}\sum_{\begin{subarray}{c}l\sim L\\ (l,fk)=1\end{subarray}}\lambda_{l}+O\left(\frac{\left(\sum_{l\sim L}|\lambda_{l}|^{2}\right)^{\frac{1}{2}}L^{\frac{1}{2}}(d(f))^{B}}{(\log L)^{A}}\right).$
(Condition 3.2: No small prime factors) We have $\lambda_{l}=0$ whenever $l$ has a prime factor less than $\exp\left(\log x(\log\log x)^{-2}\right)$.
(Condition 3.3: Values often larger than $1$) We have
$L^{1-\varepsilon}\sum_{l\sim L}|\lambda_{l}|^{4}\ll\left(\sum_{l\sim L}|\lambda_{l}|^{2}\right)^{2}.$
Lemma 3.1.
([[1], Lemma 3], [[16], Lemma 8.1]). Let $M_{1}M_{2}=x$ and $\min(M_{1},M_{2})>x^{\varepsilon^{2}}$. Suppose that $a_{2,m_{2}}$ satisfies Conditions 3.1–$3.$ If we have
$x^{\varepsilon^{2}-1}Q^{2}<M_{2}<x^{\frac{5}{6}-\varepsilon^{2}}Q^{-\frac{4}{3}},$
then
$\sum_{q\sim Q}\gamma_{q}\sum_{\begin{subarray}{c}m_{1}m_{2}\in\mathcal{A}^{q}\\ m_{i}\sim M_{i}\\ 1\leqslant i\leqslant 2\end{subarray}}a_{1,m_{1}}a_{2,m_{2}}=(1+o(1))\frac{1}{\varphi(q)}\sum_{q\sim Q}\gamma_{q}\sum_{\begin{subarray}{c}m_{1}m_{2}\in\mathcal{B}^{q}\\ m_{i}\sim M_{i}\\ 1\leqslant i\leqslant 2\end{subarray}}a_{1,m_{1}}a_{2,m_{2}}.$
Proof.
This follows from [[5], Theorem 3]. ∎
Lemma 3.2.
([[1], Lemma 4], [[16], Lemma 8.2]). Let $M_{1}M_{2}M_{3}=x$, $\min(M_{1},M_{2},M_{3})>x^{\varepsilon^{2}}$ and $A>0$. Suppose that $a_{j,m_{j}}$ ($j=1,2,3$) satisfy Condition 3.2 and $a_{2,m_{2}}$ also satisfies Condition 3.1. If there exists $B=B(A)$ such that the following conditions
$Q(\log x)^{B}<M_{1}M_{2},$
$M_{1}^{2}M_{2}^{3}<Qx^{1-\varepsilon^{2}},$
$M_{1}^{5}M_{2}^{2}<x^{2-\varepsilon^{2}}$
and
$M_{1}^{4}M_{2}^{3}<x^{2-\varepsilon^{2}}$
hold, then
$\sum_{q\sim Q}\gamma_{q}\sum_{\begin{subarray}{c}m_{1}m_{2}m_{3}\in\mathcal{A}^{q}\\ m_{i}\sim M_{i}\\ 1\leqslant i\leqslant 3\end{subarray}}a_{1,m_{1}}a_{2,m_{2}}a_{3,m_{3}}=(1+o(1))\frac{1}{\varphi(q)}\sum_{q\sim Q}\gamma_{q}\sum_{\begin{subarray}{c}m_{1}m_{2}m_{3}\in\mathcal{B}^{q}\\ m_{i}\sim M_{i}\\ 1\leqslant i\leqslant 3\end{subarray}}a_{1,m_{1}}a_{2,m_{2}}a_{3,m_{3}}.$
Proof.
This follows from [[6], Theorem 3]. ∎
Lemma 3.3.
([[1], Lemma 5], [[16], Lemma 8.3]). Let $M_{1}M_{2}M_{3}=x$, $\min(M_{1},M_{2},M_{3})>x^{\varepsilon^{2}}$ and $A>0$. Suppose that $a_{j,m_{j}}$ ($j=1,2,3$) satisfy Condition 3.2 and $a_{2,m_{2}}$ also satisfies Condition 3.1. If there exists $B=B(A)$ such that the following conditions
$Q(\log x)^{B}<M_{1}M_{2},$
$M_{1}M_{2}^{2}Q^{2}<x^{2-3\varepsilon^{2}}$
and
$M_{1}^{5}M_{2}^{2}<x^{2-3\varepsilon^{2}}$
hold, then
$\sum_{q\sim Q}\gamma_{q}\sum_{\begin{subarray}{c}m_{1}m_{2}m_{3}\in\mathcal{A}^{q}\\ m_{i}\sim M_{i}\\ 1\leqslant i\leqslant 3\end{subarray}}a_{1,m_{1}}a_{2,m_{2}}a_{3,m_{3}}=(1+o(1))\frac{1}{\varphi(q)}\sum_{q\sim Q}\gamma_{q}\sum_{\begin{subarray}{c}m_{1}m_{2}m_{3}\in\mathcal{B}^{q}\\ m_{i}\sim M_{i}\\ 1\leqslant i\leqslant 3\end{subarray}}a_{1,m_{1}}a_{2,m_{2}}a_{3,m_{3}}.$
Proof.
This is a slight variant of [[6], Theorem 4]. ∎
Lemma 3.4.
Let $M_{1}M_{2}M_{3}=x$, $\min(M_{1},M_{2},M_{3})>x^{\varepsilon^{2}}$ and $Q<x^{0.7-\varepsilon^{2}}$. Suppose that $a_{j,m_{j}}$ ($j=1,2,3$) satisfy Condition 3.2 and $a_{2,m_{2}}$ also satisfies Condition 3.1. If the following conditions
$Qx^{\varepsilon^{2}}<M_{1}M_{2},$
$M_{2}<Q^{-1}x^{1-2\varepsilon^{2}},$
$M_{1}M_{2}<Q^{-\frac{1}{7}}x^{\frac{153}{224}-10\varepsilon^{2}}$
and
$M_{1}^{4}M_{2}<Q^{-1}x^{\frac{57}{32}-10\varepsilon^{2}}$
hold, then
$\sum_{q\sim Q}\gamma_{q}\sum_{\begin{subarray}{c}m_{1}m_{2}m_{3}\in\mathcal{A}^{q}\\ m_{i}\sim M_{i}\\ 1\leqslant i\leqslant 3\end{subarray}}a_{1,m_{1}}a_{2,m_{2}}a_{3,m_{3}}=(1+o(1))\frac{1}{\varphi(q)}\sum_{q\sim Q}\gamma_{q}\sum_{\begin{subarray}{c}m_{1}m_{2}m_{3}\in\mathcal{B}^{q}\\ m_{i}\sim M_{i}\\ 1\leqslant i\leqslant 3\end{subarray}}a_{1,m_{1}}a_{2,m_{2}}a_{3,m_{3}}.$
Proof.
This follows from [[28], Proposition 8.3]. ∎
4. Sieve asymtotic formulas
In this section we give asymptotic formulas for sums of sieve functions $S\left(\mathcal{A}_{p_{1}\cdots p_{n}},p_{n}\right)$ and $S\left(\mathcal{A}_{p_{1}\cdots p_{n}},x^{\kappa_{0}}\right)$ with $\kappa_{0}=\kappa$ or $\kappa^{\prime}$, where
[…]
and
[…]
We also write
[…]
and
[…]
Now, we put $p_{i}=x^{\alpha_{i}}$ and write $\boldsymbol{\alpha}_{n}$ to denote $(\alpha_{1},\ldots,\alpha_{n})$. We shall use the terms partition and exactly partition many times in the rest of our paper. One can see [[16], Page 162] for the definition of them. Next, we shall define some regions that represent parts of the conditions of the asymptotic formulas. Note that all regions below are defined on $\theta\in[\frac{1}{2},\frac{4}{7}]$.
$\boldsymbol{g}_{1}=$ $\ \left\{(s,t):2\theta-1<s<\frac{5-8\theta}{6}\right\},$
$\boldsymbol{g}_{2}=$ $\ \left\{(s,t):s+t>\theta,\ 2s+3t<1+\theta,\ 5s+2t<2,\ 4s+3t<2\right\},$
$\boldsymbol{g}_{3}=$ $\ \left\{(s,t):s+t>\theta,\ s+2t<2-2\theta,\ 5s+2t<2\right\},$
$\boldsymbol{g}_{4}=$ $\ \left\{(s,t):s+t>\theta,\ t<1-\theta,\ s+t<\frac{153}{224}-\frac{1}{7}\theta,\ 4s+t<\frac{57}{32}-\theta\right\},$
$\boldsymbol{A}_{j}=$ $\ \left\{\boldsymbol{\alpha}_{j}:\varepsilon^{2}\leqslant\alpha_{j}<\cdots<\alpha_{1}<\tau,\ \alpha_{1}+\cdots+\alpha_{j}\leqslant 1\right\},$
$\boldsymbol{T}^{*}=$ $\ \left\{(s,t):0\leqslant s\leqslant\frac{8\theta-2}{7},\ 0\leqslant t\leqslant\frac{5-6\theta}{7}\right\},$
$\boldsymbol{T}_{j}^{*}=$ $\ \left\{\boldsymbol{\alpha}_{j}:\boldsymbol{\alpha}_{j}\text{ partitions exactly into }\boldsymbol{T}^{*}\right\},$
$\boldsymbol{S}=$ $\ \left\{(s,t):s<1-\theta,\ s+2t<2-2\theta,\ s+4t<2-\theta\right\},$
$\boldsymbol{S}_{j}=$ $\ \left\{\boldsymbol{\alpha}_{j}:\boldsymbol{\alpha}_{j}\text{ partitions exactly into }\boldsymbol{S}\right\},$
$\boldsymbol{U}_{j}^{\prime}=\boldsymbol{U}_{j}^{\prime}(\theta)=$ $\ \left\{\boldsymbol{\alpha}_{j}:\boldsymbol{\alpha}_{j}\in\boldsymbol{A}_{j},\ (\alpha_{1},\ldots,\alpha_{n},2\theta-1+\varepsilon)\in\boldsymbol{S}_{j+1}\right\},$
[…]
$\boldsymbol{T}^{**}=$ $\ \left\{(s,t):\frac{3}{7}<s<1-\theta,\ 0\leqslant t\leqslant\frac{1-s}{2}\right\},$
$\boldsymbol{U}_{j}^{*}=$ $\ \left\{\boldsymbol{\alpha}_{j}:\boldsymbol{\alpha}_{j}\text{ partitions exactly into }\boldsymbol{T}^{**}\right\},$
$\boldsymbol{G}_{j}=$ $\ \left\{\boldsymbol{\alpha}_{j}:\boldsymbol{\alpha}_{j}\in\boldsymbol{U}_{j}^{*},\text{ or }\boldsymbol{\alpha}_{j}\text{ partitions exactly into }\boldsymbol{g}_{1}\cup\boldsymbol{g}_{2}\cup\boldsymbol{g}_{3}\cup\boldsymbol{g}_{4}\right\},$
$\boldsymbol{R}=$ $\ \left\{\boldsymbol{\alpha}_{2}:\alpha_{2}\leqslant\alpha_{1},\ \alpha_{1}+2\alpha_{2}\leqslant 1,\ \alpha_{1}+4\alpha_{2}\geqslant 3-3\theta,\ 3\alpha_{2}\geqslant 2\alpha_{1},\right$.
$\left.\quad\max\left(\tau,\ \frac{31\theta-15}{3}\right)\leqslant\alpha_{1}\leqslant\min\left(\frac{3}{7},\ 4-7\theta\right)\right\},$
$\boldsymbol{R}_{0}=$ $\ \left\{\boldsymbol{\alpha}_{2}:\alpha_{1}+2\alpha_{2}\leqslant 1,\ \alpha_{1}+4\alpha_{2}\geqslant 3-3\theta,\ 3\alpha_{2}\geqslant 2\alpha_{1},\right$.
$\left.\quad\max\left(\frac{19\theta-7}{7},\ \frac{50\theta-19}{17}\right)\leqslant\alpha_{1}\leqslant\frac{3}{7}\right\},$
$\boldsymbol{D}_{1}=$ $\ \left\{\boldsymbol{\alpha}_{3}:\alpha_{1}\geqslant\kappa,\ \alpha_{2}\geqslant\kappa,\ \alpha_{3}\geqslant\kappa,\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\left.\quad(\alpha_{1},\alpha_{2}+\alpha_{3})\in\boldsymbol{R}\text{ with }\alpha_{2}\geqslant\alpha_{3}\text{ or }(\alpha_{1}+\alpha_{2},\alpha_{3})\in\boldsymbol{R}\text{ with }\alpha_{1}\geqslant\alpha_{2}\right\},$
$\boldsymbol{D}_{2}=$ $\ \left\{\boldsymbol{\alpha}_{4}:\alpha_{1}\geqslant\kappa,\ \alpha_{2}\geqslant\kappa,\ \alpha_{3}\geqslant\kappa,\ \alpha_{4}\geqslant\kappa,\right$.
$\left.\quad\boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},\ (\alpha_{1}+\alpha_{2},\alpha_{3}+\alpha_{4})\in\boldsymbol{R}\right\},$
$\boldsymbol{D}_{3}=$ $\ \left\{\boldsymbol{\alpha}_{3}:\alpha_{1}\geqslant\kappa,\ \alpha_{2}\geqslant\kappa,\ \alpha_{3}\geqslant\kappa,\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\left.\quad(\alpha_{1},\alpha_{2}+\alpha_{3})\in\boldsymbol{R}_{0}\text{ with }\alpha_{2}\geqslant\alpha_{3}\text{ or }(\alpha_{1}+\alpha_{2},\alpha_{3})\in\boldsymbol{R}_{0}\text{ with }\alpha_{1}\geqslant\alpha_{2}\right\},$
$\boldsymbol{D}_{4}=$ $\ \left\{\boldsymbol{\alpha}_{4}:\alpha_{1}\geqslant\kappa,\ \alpha_{2}\geqslant\kappa,\ \alpha_{3}\geqslant\kappa,\ \alpha_{4}\geqslant\kappa,\right$.
$\left.\quad\boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},\ (\alpha_{1}+\alpha_{2},\alpha_{3}+\alpha_{4})\in\boldsymbol{R}_{0}\right\},$
$U_{2}=$ $\ \left\{\boldsymbol{\alpha}_{2}:\frac{1}{2}\leqslant\theta<\frac{17}{32},\ \kappa\leqslant\alpha_{1}\leqslant\frac{3}{7},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$A=$ $\ \left\{\boldsymbol{\alpha}_{2}:\boldsymbol{\alpha}_{2}\in U_{2},\ \boldsymbol{\alpha}_{2}\notin\boldsymbol{G}_{2},\ \alpha_{1}+\alpha_{2}<\theta\right\},$
$B=$ $\ \left\{\boldsymbol{\alpha}_{2}:\boldsymbol{\alpha}_{2}\in U_{2},\ \boldsymbol{\alpha}_{2}\notin\boldsymbol{G}_{2}\cup A,\ \alpha_{1}+4\alpha_{2}<3-3\theta\right\},$
$C=$ $\ \left\{\boldsymbol{\alpha}_{2}:\boldsymbol{\alpha}_{2}\in U_{2},\ \boldsymbol{\alpha}_{2}\notin\boldsymbol{G}_{2}\cup A\cup B\right\}.$
Then we list some asymptotic formulas for sums whose variables lie in some suitable intervals. Note that from here we write $\mathcal{A}^{q}$ and $\mathcal{B}^{q}$ as $\mathcal{A}$ and $\mathcal{B}$ for simplicity, and the readers should remember the outer summation over $q$.
Lemma 4.1.
([[1], Lemmas 7 and 18], [[16], Lemmas 8.6 and 8.14]). Let $\frac{1}{2}\leqslant\theta<\frac{4}{7}$ and suppose that $\min\boldsymbol{\alpha}_{j}\geqslant\varepsilon^{2}$. Then we have, for almost all $q\sim x^{\theta}$,
$\sum_{\boldsymbol{\alpha}_{j}\in\boldsymbol{G}_{j}}S\left(\mathcal{A}_{p_{1}\cdots p_{j}},p_{j}\right)=(1+o(1))\frac{1}{\varphi(q)}\sum_{\boldsymbol{\alpha}_{j}\in\boldsymbol{G}_{j}}S\left(\mathcal{B}_{p_{1}\cdots p_{j}},p_{j}\right).$
Lemma 4.2.
([[1], Lemmas 15 and 16], [[16], Lemmas 8.11 and 8.12]). Let $\frac{1}{2}\leqslant\theta<\frac{4}{7}$. Then we have, for almost all $q\sim x^{\theta}$,
$\sum_{\boldsymbol{\alpha}_{j}\in\boldsymbol{U}_{j}}S\left(\mathcal{A}_{p_{1}\cdots p_{j}},x^{\kappa}\right)=(1+o(1))\frac{1}{\varphi(q)}\sum_{\boldsymbol{\alpha}_{j}\in\boldsymbol{U}_{j}}S\left(\mathcal{B}_{p_{1}\cdots p_{j}},x^{\kappa}\right).$
Let $\frac{1}{2}\leqslant\theta<\frac{11}{20}$. Then we have, for almost all $q\sim x^{\theta}$,
$\sum_{\boldsymbol{\alpha}_{j}\in\boldsymbol{U}_{j}^{\prime}}S\left(\mathcal{A}_{p_{1}\cdots p_{j}},x^{\kappa^{\prime}}\right)=(1+o(1))\frac{1}{\varphi(q)}\sum_{\boldsymbol{\alpha}_{j}\in\boldsymbol{U}_{j}^{\prime}}S\left(\mathcal{B}_{p_{1}\cdots p_{j}},x^{\kappa^{\prime}}\right).$
Lemma 4.3.
([[16], Lemma 8.15]). Let $\frac{1}{2}\leqslant\theta<\frac{17}{32}$ and $M<x^{2-\varepsilon}Q^{-3}$. Then we have, for almost all $q\sim x^{\theta}$ and a divisor-bounded sequence $a_{m}$,
$\sum_{m\sim M}a_{m}S\left(\mathcal{A}_{m},x^{\kappa}\right)=(1+o(1))\frac{1}{\varphi(q)}\sum_{m\sim M}a_{m}S\left(\mathcal{B}_{m},x^{\kappa}\right).$
Note that we have $x^{\frac{3}{7}}<x^{2-\varepsilon}Q^{-3}$ when $\theta<\frac{11}{21}-\varepsilon$.
Lemma 4.4.
([[16], Lemma 8.16]). Let $\frac{1}{2}\leqslant\theta<\frac{11}{21}$. Then we have, for almost all $q\sim x^{\theta}$,
$\sum_{\boldsymbol{\alpha}_{2}\in A\cup B}S\left(\mathcal{A}_{p_{1}p_{2}},x^{\kappa}\right)=(1+o(1))\frac{1}{\varphi(q)}\sum_{\boldsymbol{\alpha}_{2}\in A\cup B}S\left(\mathcal{B}_{p_{1}p_{2}},x^{\kappa}\right).$
5. High-dimensional sieves
In this section, we mention several results regarding the upper bounds for some sieve functions. These upper bounds are proved using two- or three-dimensional sieves.
Lemma 5.1.
([[1], Lemma 20], [[16], Lemma 8.18]). Let $\frac{1}{2}\leqslant\theta<\frac{17}{32}$. Then we have, for almost all $q\sim x^{\theta}$,
$\sum_{1-\theta\leqslant\alpha_{1}<\frac{1}{2}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)\leqslant(1+o(1))\frac{1}{\varphi(q)}\left(-\sum_{1-\theta\leqslant\alpha_{1}<\frac{1}{2}}S\left(\mathcal{B}_{p_{1}},p_{1}\right)+\sum_{\begin{subarray}{c}\kappa\leqslant\alpha_{3}<\alpha_{1}<\frac{\theta}{2}\\ 1-\theta\leqslant\alpha_{1}+\beta\leqslant\theta\\ \alpha_{3}<\min\left(\beta,\frac{1}{2}(1-\alpha_{1}-\beta)\right)\\ (\alpha_{1},\beta,\alpha_{3})\notin\boldsymbol{G}_{3}\end{subarray}}S\left(\mathcal{B}_{p_{1}mp_{3}},p_{3}\right)\right),$
where $m=x^{\beta}$ and $\left(m,P(p_{1})\right)=1$.
Lemma 5.2.
([[1], Lemma 24], [[16], Lemma 8.23]). Let $\frac{1}{2}\leqslant\theta<\frac{16}{31}$. Then we have, for almost all $q\sim x^{\theta}$,
$\sum_{\boldsymbol{\alpha}_{2}\in\boldsymbol{R}}S\left(\mathcal{A}_{p_{1}p_{2}},x^{\kappa}\right)\leqslant(1+o(1))\frac{1}{\varphi(q)}\left(\sum_{\boldsymbol{\alpha}_{2}\in\boldsymbol{R}}S\left(\mathcal{B}_{p_{1}p_{2}},x^{\kappa}\right)+\frac{x}{\log x}(I_{1}+I_{2})\right),$
where
$I_{1}=\frac{1}{\kappa}\int_{(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{1}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1},$
$I_{2}=\frac{1}{\kappa}\int_{(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{2}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1},$
where $\omega(u)$, here and below, denote the Buchstab function determined by the following differential–difference equation
$\begin{cases}\omega(u)=\frac{1}{u},&\quad 1\leqslant u\leqslant 2,\\ (u\omega(u))^{\prime}=\omega(u-1),&\quad u\geqslant 2.\end{cases}$
Lemma 5.3.
([[2], Lemma 24]). Let $\frac{25}{49}\leqslant\theta\leqslant\frac{92}{175}$. Then we have, for almost all $q\sim x^{\theta}$,
$\sum_{\boldsymbol{\alpha}_{2}\in\boldsymbol{R}_{0}}S\left(\mathcal{A}_{p_{1}p_{2}},x^{\kappa}\right)\leqslant(1+o(1))\frac{1}{\varphi(q)}\left(\sum_{\boldsymbol{\alpha}_{2}\in\boldsymbol{R}_{0}}S\left(\mathcal{B}_{p_{1}p_{2}},x^{\kappa}\right)+\frac{x}{\log x}(I_{3}+I_{4})\right),$
where
$I_{3}=\frac{1}{\kappa}\int_{(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{3}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1},$
$I_{4}=\frac{1}{\kappa}\int_{(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{4}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}.$
6. The final decomposition
In this section, we ignore the presence of $\varepsilon$ for clarity. Let $\omega(u)$ denote the Buchstab function defined in Lemma 5.2. Let $\frac{1}{2}\leqslant\theta\leqslant 0.679$. We shall split this range of $\theta$ to several subranges and use different methods to treat them and obtain good bounds for $C(\theta)$.
6.1. Case 1. $\frac{1}{2}\leqslant\theta<\frac{25}{49}$
In the first subrange, we can use the lemmas in Sections 4 and 5 with their full power. Using Buchstab’s identity, we have
$S\left(\mathcal{A},(2x)^{\frac{1}{2}}\right)=$ $\ S\left(\mathcal{A},x^{\kappa}\right)-\sum_{\kappa\leqslant\alpha_{1}<\frac{1}{2}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)$
$=$ $\ S\left(\mathcal{A},x^{\kappa}\right)-\sum_{\kappa\leqslant\alpha_{1}\leqslant\frac{3}{7}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)-\sum_{\frac{3}{7}<\alpha_{1}<1-\theta}S\left(\mathcal{A}_{p_{1}},p_{1}\right)-\sum_{1-\theta\leqslant\alpha_{1}<\frac{1}{2}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)$
$=$ $\ S_{11}-S_{12}-S_{13}-S_{14}.$ (17)
By Lemma 4.3, we can give an asymptotic formula for $S_{11}$. By Lemma 4.1, we can also give an asymptotic formula for $S_{13}$. For $S_{14}$, we can use Lemma 5.1 to give an upper bound with a loss of
$\sum_{\begin{subarray}{c}\kappa\leqslant\alpha_{3}<\alpha_{1}<\frac{\theta}{2}\\ 1-\theta\leqslant\alpha_{1}+\beta\leqslant\theta\\ \alpha_{3}<\min\left(\beta,\frac{1}{2}(1-\alpha_{1}-\beta)\right)\\ (\alpha_{1},\beta,\alpha_{3})\notin\boldsymbol{G}_{3}\end{subarray}}S\left(\mathcal{A}_{p_{1}mp_{3}},p_{3}\right).$ (18)
For the remaining $S_{12}$, we perform Buchstab’s identity twice again to get
$S_{12}=$ $\ \sum_{\kappa\leqslant\alpha_{1}\leqslant\frac{3}{7}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)$
$=$ $\ \sum_{\kappa\leqslant\alpha_{1}\leqslant\frac{3}{7}}S\left(\mathcal{A}_{p_{1}},x^{\kappa}\right)-\sum_{\begin{subarray}{c}\kappa\leqslant\alpha_{1}\leqslant\frac{3}{7}\\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}},p_{2}\right)$
$=$ $\ \sum_{\kappa\leqslant\alpha_{1}\leqslant\frac{3}{7}}S\left(\mathcal{A}_{p_{1}},x^{\kappa}\right)-\sum_{\begin{subarray}{c}\kappa\leqslant\alpha_{1}\leqslant\frac{3}{7}\\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\\ \boldsymbol{\alpha}_{2}\in\boldsymbol{G}_{2}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}},p_{2}\right)-\sum_{\begin{subarray}{c}\kappa\leqslant\alpha_{1}\leqslant\frac{3}{7}\\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\\ \boldsymbol{\alpha}_{2}\notin\boldsymbol{G}_{2}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}},p_{2}\right)$
$=$ $\ S_{121}-S_{122}-S_{123}.$ (19)
We can give asymptotic formulas for $S_{121}$ by Lemma 4.3 and for $S_{122}$ by Lemma 4.1. For $S_{123}$, we can easily find that $\boldsymbol{\alpha}_{2}\in U_{2}$. Thus, we split it into two parts $\boldsymbol{\alpha}_{2}\in A\cup B$ and $\boldsymbol{\alpha}_{2}\in C$:
$S_{123}=$ $\ \sum_{\boldsymbol{\alpha}_{2}\in A\cup B}S\left(\mathcal{A}_{p_{1}p_{2}},p_{2}\right)+\sum_{\boldsymbol{\alpha}_{2}\in C}S\left(\mathcal{A}_{p_{1}p_{2}},p_{2}\right)$
$=$ $\ \sum_{\boldsymbol{\alpha}_{2}\in A\cup B}S\left(\mathcal{A}_{p_{1}p_{2}},x^{\kappa}\right)-\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{2}\in A\cup B\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\in\boldsymbol{G}_{3}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}},p_{3}\right)-\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{2}\in A\cup B\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}},p_{3}\right)$
$+\sum_{\boldsymbol{\alpha}_{2}\in C}S\left(\mathcal{A}_{p_{1}p_{2}},x^{\kappa}\right)-\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{2}\in C\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\in\boldsymbol{G}_{3}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}},p_{3}\right)-\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{2}\in C\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}},p_{3}\right)$
$=$ $\ S_{1231}-S_{1232}-S_{1233}+S_{1234}-S_{1235}-S_{1236}.$ (20)
For $S_{1231}$, we can use Lemma 4.4 to give an asymptotic formula. We can also give asymptotic formulas for $S_{1232}$ and $S_{1235}$ by Lemma 4.1. However, for $S_{1234}$ we cannot give an asymptotic formula. Note that we cannot discard any part of $S_{1234}$ since it has a positive sign. Now we must use Lemma 5.2 to give an upper bound for this sum. After checking the boundaries of the region $\boldsymbol{R}$, we find that $\boldsymbol{R}$ covers the whole region $C$, and the loss from $S_{1234}$ is the sum of two corresponding integrals $I_{1}$ and $I_{2}$. We discard the whole sum $S_{1236}$.
Now we only need to deal with the remaining sum $S_{1233}$. We want to use Buchstab’s identity twice again, hence we need to check whether the conditions in Lemma 4.2 are fulfilled. That is, if we have $\boldsymbol{\alpha}_{3}\in\boldsymbol{U}_{3}$ and
$(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\in\boldsymbol{U}_{4},$
for parts of $S_{1233}$ (say, $S_{1233}^{\prime}$), then we can use Buchstab’s identity twice again to get
$S_{1233}^{\prime}=$ $\ \sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{2}\in A\cup B\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\\ \boldsymbol{\alpha}_{3}\in\boldsymbol{U}_{3}\\ (\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\in\boldsymbol{U}_{4}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}},p_{3}\right)$
$=$ $\ \sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{2}\in A\cup B\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\boldsymbol{\alpha}_{3}\in\boldsymbol{U}_{3}\\ (\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\in\boldsymbol{U}_{4}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}},x^{\kappa}\right)-\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{2}\in A\cup B\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\\ \boldsymbol{\alpha}_{3}\in\boldsymbol{U}_{3}\\ (\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\in\boldsymbol{U}_{4}\\ \kappa\leqslant\alpha_{4}<\min\left(\alpha_{3},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\right)\\ \boldsymbol{\alpha}_{4}\in\boldsymbol{G}_{4}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}p_{4}},p_{4}\right)$
$-\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{2}\in A\cup B\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\\ \boldsymbol{\alpha}_{3}\in\boldsymbol{U}_{3}\\ (\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\in\boldsymbol{U}_{4}\\ \kappa\leqslant\alpha_{4}<\min\left(\alpha_{3},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\right)\\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}p_{4}},x^{\kappa}\right)+\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{2}\in A\cup B\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\\ \boldsymbol{\alpha}_{3}\in\boldsymbol{U}_{3}\\ (\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\in\boldsymbol{U}_{4}\\ \kappa\leqslant\alpha_{4}<\min\left(\alpha_{3},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\right)\\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4}\\ \kappa\leqslant\alpha_{5}<\min\left(\alpha_{4},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4})\right)\\ \boldsymbol{\alpha}_{5}\in\boldsymbol{G}_{5}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}p_{4}p_{5}},p_{5}\right)$
$+\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{2}\in A\cup B\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\\ \boldsymbol{\alpha}_{3}\in\boldsymbol{U}_{3}\\ (\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\in\boldsymbol{U}_{4}\\ \kappa\leqslant\alpha_{4}<\min\left(\alpha_{3},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\right)\\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4}\\ \kappa\leqslant\alpha_{5}<\min\left(\alpha_{4},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4})\right)\\ \boldsymbol{\alpha}_{5}\notin\boldsymbol{G}_{5}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}p_{4}p_{5}},p_{5}\right)$
$=$ $\ S_{12331}-S_{12332}-S_{12333}+S_{12334}+S_{12335}.$ (21)
We can give asymptotic formulas for $S_{12332}$ and $S_{12334}$ by Lemma 4.1. Since $\boldsymbol{\alpha}_{3}\in\boldsymbol{U}_{3}$, we can use Lemma 4.2 to give an asymptotic formula for $S_{12331}$. Again, since $\alpha_{4}<\alpha_{3}$ and $(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\in\boldsymbol{U}_{4}$, we have $\boldsymbol{\alpha}_{4}\in\boldsymbol{U}_{4}$ and thus Lemma 4.2 is usable for $S_{12333}$. We discard $S_{12335}$ and the remaining parts of $S_{1233}$, leading to a three-dimensional integral and a five-dimensional integral of loss. Note that we can use the reversed Buchstab’s identity to gain a small saving over the remaining parts of $S_{1233}$. The process can be seen as the following:
$S_{1233}-S_{1233}^{\prime}=$ $\ \sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{2}\in A\cup B\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\\ \text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}},p_{3}\right)$
$=$ $\ \sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{2}\in A\cup B\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\\ \text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4}\\ \alpha_{3}\geqslant\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}},p_{3}\right)+\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{2}\in A\cup B\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\\ \text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4}\\ \alpha_{3}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}},p_{3}\right)$
$=$ $\ \sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{2}\in A\cup B\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\\ \text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4}\\ \alpha_{3}\geqslant\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}},p_{3}\right)+\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{2}\in A\cup B\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\\ \text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4}\\ \alpha_{3}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}},\left(\frac{x}{p_{1}p_{2}p_{3}}\right)^{\frac{1}{2}}\right)$
$+\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{2}\in A\cup B\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\\ \text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4}\\ \alpha_{3}<\alpha_{4}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\\ \boldsymbol{\alpha}_{4}\in\boldsymbol{G}_{4}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}p_{4}},p_{4}\right)+\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{2}\in A\cup B\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\\ \text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4}\\ \alpha_{3}<\alpha_{4}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4}\\ \alpha_{4}\geqslant\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4})\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}p_{4}},p_{4}\right)$
$+\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{2}\in A\cup B\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\\ \text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4}\\ \alpha_{3}<\alpha_{4}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4}\\ \alpha_{4}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4})\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}p_{4}},p_{4}\right)$
$=$ $\ \sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{2}\in A\cup B\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\\ \text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4}\\ \alpha_{3}\geqslant\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}},p_{3}\right)+\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{2}\in A\cup B\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\\ \text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4}\\ \alpha_{3}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}},\left(\frac{x}{p_{1}p_{2}p_{3}}\right)^{\frac{1}{2}}\right)$
$+\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{2}\in A\cup B\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\\ \text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4}\\ \alpha_{3}<\alpha_{4}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\\ \boldsymbol{\alpha}_{4}\in\boldsymbol{G}_{4}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}p_{4}},p_{4}\right)+\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{2}\in A\cup B\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\\ \text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4}\\ \alpha_{3}<\alpha_{4}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4}\\ \alpha_{4}\geqslant\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4})\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}p_{4}},p_{4}\right)$
$+\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{2}\in A\cup B\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\\ \text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4}\\ \alpha_{3}<\alpha_{4}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4}\\ \alpha_{4}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4})\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}p_{4}},\left(\frac{x}{p_{1}p_{2}p_{3}p_{4}}\right)^{\frac{1}{2}}\right)$
$+\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{2}\in A\cup B\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\\ \text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4}\\ \alpha_{3}<\alpha_{4}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4}\\ \alpha_{4}<\alpha_{5}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4})\\ \boldsymbol{\alpha}_{5}\in\boldsymbol{G}_{5}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}p_{4}p_{5}},p_{5}\right)$
$+\sum_{\begin{subarray}{c}\boldsymbol{\alpha}_{2}\in A\cup B\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\\ \text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4}\\ \alpha_{3}<\alpha_{4}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4}\\ \alpha_{4}<\alpha_{5}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4})\\ \boldsymbol{\alpha}_{5}\notin\boldsymbol{G}_{5}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}p_{4}p_{5}},p_{5}\right)$
$=$ $\ S_{1241}+S_{1242}+S_{1243}+S_{1244}+S_{1245}+S_{1246}+S_{1247}.$ (22)
We can use Lemma 4.1 to deal with $S_{1243}$ and $S_{1246}$. Therefore, we can subtract them from the loss of $S_{1233}$. By (17)–(22), we get the total loss when $\frac{1}{2}\leqslant\theta<\frac{25}{49}$ of
$\left(\int_{\kappa}^{\frac{\theta}{2}}\int_{1-\theta-t_{1}}^{\theta-t_{1}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{101}]\frac{\omega\left(\frac{t_{2}}{t_{1}}\right)\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}^{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{1}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{2}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{104}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{105}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in L_{106}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\int_{t_{4}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{107}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{108}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$=$ $\ E_{101}(\theta)+E_{102}(\theta)+E_{103}(\theta)+E_{104}(\theta)+E_{105}(\theta)-E_{106}(\theta)-E_{107}(\theta)+E_{108}(\theta),$ (23)
where
$L_{101}(\boldsymbol{\alpha}_{3}):=$ $\ \left\{\kappa\leqslant\alpha_{1}\leqslant\frac{\theta}{2},\ 1-\theta-\alpha_{1}\leqslant\alpha_{2}\leqslant\theta-\alpha_{1},\right$.
$\left.\quad\kappa\leqslant\alpha_{3}\leqslant\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\right\},$
$L_{104}(\boldsymbol{\alpha}_{3}):=$ $\ \left\{\boldsymbol{\alpha}_{2}\in C,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{105}(\boldsymbol{\alpha}_{3}):=$ $\ \left\{\boldsymbol{\alpha}_{2}\in A\cup B,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{106}(\boldsymbol{\alpha}_{4}):=$ $\ \left\{\boldsymbol{\alpha}_{2}\in A\cup B,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4},$
$\quad\alpha_{3}<\alpha_{4}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}),\ \boldsymbol{\alpha}_{4}\in\boldsymbol{G}_{4},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{107}(\boldsymbol{\alpha}_{5}):=$ $\ \left\{\boldsymbol{\alpha}_{2}\in A\cup B,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4},$
$\quad\alpha_{3}<\alpha_{4}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}),\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},$
$\quad\alpha_{4}<\alpha_{5}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4}),\ \boldsymbol{\alpha}_{5}\in\boldsymbol{G}_{5},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{108}(\boldsymbol{\alpha}_{5}):=$ $\ \left\{\boldsymbol{\alpha}_{2}\in A\cup B,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\boldsymbol{\alpha}_{3}\in\boldsymbol{U}_{3},\ (\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\in\boldsymbol{U}_{4},$
$\quad\kappa\leqslant\alpha_{4}<\min\left(\alpha_{3},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\right),\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},$
$\quad\kappa\leqslant\alpha_{5}<\min\left(\alpha_{4},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4})\right),\ \boldsymbol{\alpha}_{5}\notin\boldsymbol{G}_{5},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\}.$
Finally, for $\frac{1}{2}\leqslant\theta<\frac{25}{49}$ we have
$C(\theta)\leqslant C_{1}(\theta):=1+\left(E_{101}(\theta)+E_{102}(\theta)+E_{103}(\theta)+E_{104}(\theta)+E_{105}(\theta)-E_{106}(\theta)-E_{107}(\theta)+E_{108}(\theta)\right).$ (24)
$\theta$ $C_{1}(\theta)$
$0.501$ $1.00001$
$0.503$ $1.00124$
$0.505$ $1.00633$
$0.508$ $1.01001$
$0.51$ $1.015$
$\frac{25}{49}$ $1.02079$
6.2. Case 2. $\frac{25}{49}\leqslant\theta<\frac{21}{41}$
In this subrange, we note that $4-7\theta\leqslant\frac{3}{7}$ and the upper bound for $\alpha_{1}\in\boldsymbol{R}$ becomes $4-7\theta$. Since we cannot discard any parts of the sum
$\sum_{\boldsymbol{\alpha}_{2}\in C}S\left(\mathcal{A}_{p_{1}p_{2}},p_{2}\right),$ (25)
we must split the sum
$\sum_{\kappa\leqslant\alpha_{1}\leqslant\frac{3}{7}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)$
into two sums. Now we start our decomposition by using Buchstab’s identity:
$S\left(\mathcal{A},(2x)^{\frac{1}{2}}\right)=$ $\ S\left(\mathcal{A},x^{\kappa}\right)-\sum_{\kappa\leqslant\alpha_{1}<\frac{1}{2}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)$
$=$ $\ S\left(\mathcal{A},x^{\kappa}\right)-\sum_{\kappa\leqslant\alpha_{1}<4-7\theta}S\left(\mathcal{A}_{p_{1}},p_{1}\right)-\sum_{4-7\theta\leqslant\alpha_{1}\leqslant\frac{3}{7}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)$
$-\sum_{\frac{3}{7}<\alpha_{1}<1-\theta}S\left(\mathcal{A}_{p_{1}},p_{1}\right)-\sum_{1-\theta\leqslant\alpha_{1}<\frac{1}{2}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)$
$=$ $\ S_{21}-S_{22}-S_{23}-S_{24}-S_{25}.$ (26)
Just as $S_{11}$ and $S_{13}$ in Case 1, we can give asymptotic formulas for $S_{21}$ by Lemma 4.3 and for $S_{24}$ by Lemma 4.1. For $S_{25}$, we can use Lemma 5.1 to give an upper bound similar to that for $S_{14}$. We can use the same process as in the estimation $S_{12}$ to show a similar upper bound for $S_{22}$, with a loss consisting of 7 integrals ($I_{1}$, $I_{2}$, 3D for $\boldsymbol{\alpha}_{2}\in C$, and 3D+, 4D-, 5D- and 5D+ for $\boldsymbol{\alpha}_{2}\in A\cup B$). Now we only need to deal with the remaining $S_{23}$. In this sum Lemma 5.2 is not applicable, but we can use Lemma 5.3 to deal with some parts satisfying
$\max\left(\frac{19\theta-7}{7},\ \frac{50\theta-19}{17}\right)\leqslant\alpha_{1}\leqslant\frac{3}{7},\quad\text{or}\quad\frac{19\theta-7}{7}\leqslant\alpha_{1}\leqslant\frac{3}{7}\quad\text{in this case}.$
That is, we use Buchstab’s identity on $S_{23}$ twice again, giving the following expression.
$S_{23}=$ $\ \sum_{4-7\theta\leqslant\alpha_{1}\leqslant\frac{3}{7}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)$
$=$ $\ \sum_{\begin{subarray}{c}4-7\theta\leqslant\alpha_{1}\leqslant\frac{3}{7}\\ \theta\notin\left[\frac{19\theta-7}{7},\frac{3}{7}\right]\end{subarray}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)+\sum_{\begin{subarray}{c}4-7\theta\leqslant\alpha_{1}\leqslant\frac{3}{7}\\ \theta\in\left[\frac{19\theta-7}{7},\frac{3}{7}\right]\end{subarray}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)$
$=$ $\ \sum_{\begin{subarray}{c}4-7\theta\leqslant\alpha_{1}\leqslant\frac{3}{7}\\ \theta\notin\left[\frac{19\theta-7}{7},\frac{3}{7}\right]\end{subarray}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)+\sum_{\begin{subarray}{c}4-7\theta\leqslant\alpha_{1}\leqslant\frac{3}{7}\\ \theta\in\left[\frac{19\theta-7}{7},\frac{3}{7}\right]\end{subarray}}S\left(\mathcal{A}_{p_{1}},x^{\kappa}\right)$
$-\sum_{\begin{subarray}{c}4-7\theta\leqslant\alpha_{1}\leqslant\frac{3}{7}\\ \theta\in\left[\frac{19\theta-7}{7},\frac{3}{7}\right]\\ \boldsymbol{\alpha}_{2}\in A\cup B\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}},x^{\kappa}\right)-\sum_{\begin{subarray}{c}4-7\theta\leqslant\alpha_{1}\leqslant\frac{3}{7}\\ \theta\in\left[\frac{19\theta-7}{7},\frac{3}{7}\right]\\ \boldsymbol{\alpha}_{2}\in C\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}},x^{\kappa}\right)$
$+\sum_{\begin{subarray}{c}4-7\theta\leqslant\alpha_{1}\leqslant\frac{3}{7}\\ \theta\in\left[\frac{19\theta-7}{7},\frac{3}{7}\right]\\ \boldsymbol{\alpha}_{2}\in A\cup B\cup C\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\in\boldsymbol{G}_{3}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}},p_{3}\right)+\sum_{\begin{subarray}{c}4-7\theta\leqslant\alpha_{1}\leqslant\frac{3}{7}\\ \theta\in\left[\frac{19\theta-7}{7},\frac{3}{7}\right]\\ \boldsymbol{\alpha}_{2}\in A\cup B\cup C\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}},p_{3}\right)$
$=$ $\ S_{231}+S_{232}-S_{233}-S_{234}+S_{235}+S_{236}.$ (27)
We give asymptotic formulas for $S_{232}$ (Lemma 4.3), $S_{233}$ (Lemma 4.4) and $S_{235}$ (Lemma 4.1). After checking the conditions we find that $\boldsymbol{R}_{0}$ covers parts of $C$ satisfy $\alpha_{1}\in\left[\frac{19\theta-7}{7},\frac{3}{7}\right]$. We use Lemma 5.3 to give an upper bound for $S_{234}$ and discard the whole of $S_{231}$ and $S_{236}$, leading to 4 loss integrals ($S_{231}\to$ 1D for $4-7\theta\leqslant\alpha_{1}<\frac{19\theta-7}{7}$, $S_{234}\to$ $I_{3}$ and $I_{4}$, and $S_{236}\to$ 3D for $\frac{19\theta-7}{7}\leqslant\alpha_{1}\leqslant\frac{3}{7}$). Together, the total loss when $\frac{25}{49}\leqslant\theta<\frac{21}{41}$ can be bounded by
$\left(\int_{\kappa}^{\frac{\theta}{2}}\int_{1-\theta-t_{1}}^{\theta-t_{1}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{201}]\frac{\omega\left(\frac{t_{2}}{t_{1}}\right)\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}^{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{1}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{2}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{204}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{205}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in L_{206}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\int_{t_{4}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{207}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{208}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{4-7\theta}^{\frac{3}{7}}\texttt{Boole}[(t_{1})\in L_{209}]\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{3}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{4}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{212}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
$=$ $\ E_{201}(\theta)+E_{202}(\theta)+E_{203}(\theta)+E_{204}(\theta)+E_{205}(\theta)-E_{206}(\theta)-E_{207}(\theta)$
$+E_{208}(\theta)+E_{209}(\theta)+E_{210}(\theta)+E_{211}(\theta)+E_{212}(\theta),$ (28)
where
$L_{201}(\boldsymbol{\alpha}_{3}):=$ $\ \left\{\kappa\leqslant\alpha_{1}\leqslant\frac{\theta}{2},\ 1-\theta-\alpha_{1}\leqslant\alpha_{2}\leqslant\theta-\alpha_{1},\right$.
$\left.\quad\kappa\leqslant\alpha_{3}\leqslant\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\right\},$
$L_{204}(\boldsymbol{\alpha}_{3}):=$ $\ \left\{\alpha_{1}<4-7\theta,\ \boldsymbol{\alpha}_{2}\in C,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{205}(\boldsymbol{\alpha}_{3}):=$ $\ \left\{\alpha_{1}<4-7\theta,\ \boldsymbol{\alpha}_{2}\in A\cup B,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{206}(\boldsymbol{\alpha}_{4}):=$ $\ \left\{\alpha_{1}<4-7\theta,\ \boldsymbol{\alpha}_{2}\in A\cup B,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4},$
$\quad\alpha_{3}<\alpha_{4}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}),\ \boldsymbol{\alpha}_{4}\in\boldsymbol{G}_{4},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{207}(\boldsymbol{\alpha}_{5}):=$ $\ \left\{\alpha_{1}<4-7\theta,\ \boldsymbol{\alpha}_{2}\in A\cup B,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4},$
$\quad\alpha_{3}<\alpha_{4}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}),\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},$
$\quad\alpha_{4}<\alpha_{5}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4}),\ \boldsymbol{\alpha}_{5}\in\boldsymbol{G}_{5},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{208}(\boldsymbol{\alpha}_{5}):=$ $\ \left\{\alpha_{1}<4-7\theta,\ \boldsymbol{\alpha}_{2}\in A\cup B,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\boldsymbol{\alpha}_{3}\in\boldsymbol{U}_{3},\ (\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\in\boldsymbol{U}_{4},$
$\quad\kappa\leqslant\alpha_{4}<\min\left(\alpha_{3},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\right),\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},$
$\quad\kappa\leqslant\alpha_{5}<\min\left(\alpha_{4},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4})\right),\ \boldsymbol{\alpha}_{5}\notin\boldsymbol{G}_{5},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{209}(\alpha_{1}):=$ $\ \left\{4-7\theta\leqslant\alpha_{1}\leqslant\frac{3}{7},\ \alpha_{1}\notin\left[\frac{19\theta-7}{7},\ \frac{3}{7}\right]\right\},$
$L_{212}(\boldsymbol{\alpha}_{3}):=$ $\ \left\{4-7\theta\leqslant\alpha_{1}\leqslant\frac{3}{7},\ \alpha_{1}\in\left[\frac{19\theta-7}{7},\ \frac{3}{7}\right],\ \boldsymbol{\alpha}_{2}\notin\boldsymbol{G}_{2},\right$.
$\quad\kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\}.$
Finally, for $\frac{25}{49}\leqslant\theta<\frac{21}{41}$ we have
$C(\theta)\leqslant C_{2}(\theta):=$ $\ 1+\left(E_{201}(\theta)+E_{202}(\theta)+E_{203}(\theta)+E_{204}(\theta)+E_{205}(\theta)-E_{206}(\theta)\right$.
$\left.\qquad-E_{207}(\theta)+E_{208}(\theta)+E_{209}(\theta)+E_{210}(\theta)+E_{211}(\theta)+E_{212}(\theta)\right).$ (29)
$\theta$ $C_{2}(\theta)$ $\theta$ $C_{2}(\theta)$ $\theta$ $C_{2}(\theta)$ $\theta$ $C_{2}(\theta)$
$0.5103$ $1.02214$ $0.5109$ $1.02349$ $0.5114$ $1.02441$ $0.5119$ $1.02757$
$0.5104$ $1.02236$ $0.5111$ $1.02365$ $0.5115$ $1.02623$ $0.512$ $1.03146$
$0.5105$ $1.0227$ $0.5112$ $1.02385$ $0.51155$ $1.02639$ $0.5121$ $1.0315$
$0.5107$ $1.02297$ $0.5113$ $1.02414$ $0.5116$ $1.02656$ $\frac{21}{41}$ $1.03169$
6.3. Case 3. $\frac{21}{41}\leqslant\theta<\frac{16}{31}$
In this subrange, we have $\frac{31\theta-15}{3}\geqslant\tau$, and the lower bound for $\alpha_{1}\in\boldsymbol{R}$ becomes $\frac{31\theta-15}{3}$. Because the lines $\alpha_{1}=\alpha_{2}$ and $\alpha_{1}+4\alpha_{2}=3-3\theta$ intersect in $(\alpha_{1},\alpha_{2})=(\tau,\tau)$, we have $\boldsymbol{\alpha}_{2}\in A\cup B$ if $\alpha_{2}<\alpha_{1}<\tau$. Thus, we start our decomposition by using Buchstab’s identity and split the second sum in the following way:
$S\left(\mathcal{A},(2x)^{\frac{1}{2}}\right)=$ $\ S\left(\mathcal{A},x^{\kappa}\right)-\sum_{\kappa\leqslant\alpha_{1}<\frac{1}{2}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)$
$=$ $\ S\left(\mathcal{A},x^{\kappa}\right)-\sum_{\alpha_{1}\in\left[\kappa,\tau\right)\cup\left(\frac{31\theta-15}{3},4-7\theta\right)}S\left(\mathcal{A}_{p_{1}},p_{1}\right)-\sum_{\alpha_{1}\in\left[\tau,\frac{31\theta-15}{3}\right]\cup\left[4-7\theta,\frac{3}{7}\right]}S\left(\mathcal{A}_{p_{1}},p_{1}\right)$
$-\sum_{\frac{3}{7}<\alpha_{1}<1-\theta}S\left(\mathcal{A}_{p_{1}},p_{1}\right)-\sum_{1-\theta\leqslant\alpha_{1}<\frac{1}{2}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)$
$=$ $\ S_{31}-S_{32}-S_{33}-S_{34}-S_{35}.$ (30)
By a decomposing process similar to our decomposition in Case 2, we can get 7 loss integrals for $S_{22}$, 4 loss integrals for $S_{23}$ and a loss integral for $S_{25}$. Thus, the total loss when $\frac{21}{41}\leqslant\theta<\frac{16}{31}$ can be bounded by
$\left(\int_{\kappa}^{\frac{\theta}{2}}\int_{1-\theta-t_{1}}^{\theta-t_{1}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{301}]\frac{\omega\left(\frac{t_{2}}{t_{1}}\right)\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}^{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{1}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{2}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{304}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{305}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in L_{306}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\int_{t_{4}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{307}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{308}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\tau}^{\frac{3}{7}}\texttt{Boole}[(t_{1})\in L_{309}]\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{3}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{4}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{312}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
$=$ $\ E_{301}(\theta)+E_{302}(\theta)+E_{303}(\theta)+E_{304}(\theta)+E_{305}(\theta)-E_{306}(\theta)-E_{307}(\theta)$
$+E_{308}(\theta)+E_{309}(\theta)+E_{310}(\theta)+E_{311}(\theta)+E_{312}(\theta),$ (31)
where
$L_{301}(\boldsymbol{\alpha}_{3}):=$ $\ \left\{\kappa\leqslant\alpha_{1}\leqslant\frac{\theta}{2},\ 1-\theta-\alpha_{1}\leqslant\alpha_{2}\leqslant\theta-\alpha_{1},\right$.
$\left.\quad\kappa\leqslant\alpha_{3}\leqslant\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\right\},$
$L_{304}(\boldsymbol{\alpha}_{3}):=$ $\ \left\{\alpha_{1}\in\left[\kappa,\ \tau\right)\cup\left(\frac{31\theta-15}{3},\ 4-7\theta\right),\ \boldsymbol{\alpha}_{2}\in C,\right$.
$\quad\kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{305}(\boldsymbol{\alpha}_{3}):=$ $\ \left\{\alpha_{1}\in\left[\kappa,\ \tau\right)\cup\left(\frac{31\theta-15}{3},\ 4-7\theta\right),\ \boldsymbol{\alpha}_{2}\in A\cup B,\right$.
$\quad\kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},$
$\quad\text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{306}(\boldsymbol{\alpha}_{4}):=$ $\ \left\{\alpha_{1}\in\left[\kappa,\ \tau\right)\cup\left(\frac{31\theta-15}{3},\ 4-7\theta\right),\ \boldsymbol{\alpha}_{2}\in A\cup B,\right$.
$\quad\kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},$
$\quad\text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4},$
$\quad\alpha_{3}<\alpha_{4}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}),\ \boldsymbol{\alpha}_{4}\in\boldsymbol{G}_{4},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{307}(\boldsymbol{\alpha}_{5}):=$ $\ \left\{\alpha_{1}\in\left[\kappa,\ \tau\right)\cup\left(\frac{31\theta-15}{3},\ 4-7\theta\right),\ \boldsymbol{\alpha}_{2}\in A\cup B,\right$.
$\quad\kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},$
$\quad\text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4},$
$\quad\alpha_{3}<\alpha_{4}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}),\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},$
$\quad\alpha_{4}<\alpha_{5}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4}),\ \boldsymbol{\alpha}_{5}\in\boldsymbol{G}_{5},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{308}(\boldsymbol{\alpha}_{5}):=$ $\ \left\{\alpha_{1}\in\left[\kappa,\ \tau\right)\cup\left(\frac{31\theta-15}{3},\ 4-7\theta\right),\ \boldsymbol{\alpha}_{2}\in A\cup B,\right$.
$\quad\kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},$
$\quad\boldsymbol{\alpha}_{3}\in\boldsymbol{U}_{3},\ (\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\in\boldsymbol{U}_{4},$
$\quad\kappa\leqslant\alpha_{4}<\min\left(\alpha_{3},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\right),\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},$
$\quad\kappa\leqslant\alpha_{5}<\min\left(\alpha_{4},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4})\right),\ \boldsymbol{\alpha}_{5}\notin\boldsymbol{G}_{5},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{309}(\alpha_{1}):=$ $\ \left\{\alpha_{1}\in\left[\tau,\ \frac{31\theta-15}{3}\right]\cup\left[4-7\theta,\ \frac{3}{7}\right],\ \alpha_{1}\notin\left[\frac{19\theta-7}{7},\ \frac{3}{7}\right]\right\},$
$L_{312}(\boldsymbol{\alpha}_{3}):=$ $\ \left\{\alpha_{1}\in\left[\tau,\ \frac{31\theta-15}{3}\right]\cup\left[4-7\theta,\ \frac{3}{7}\right],\ \alpha_{1}\in\left[\frac{19\theta-7}{7},\ \frac{3}{7}\right],\ \boldsymbol{\alpha}_{2}\notin\boldsymbol{G}_{2},\right$.
$\quad\kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\}.$
Finally, for $\frac{21}{41}\leqslant\theta<\frac{16}{31}$ we have
$C(\theta)\leqslant C_{3}(\theta):=$ $\ 1+\left(E_{301}(\theta)+E_{302}(\theta)+E_{303}(\theta)+E_{304}(\theta)+E_{305}(\theta)-E_{306}(\theta)\right$.
$\left.\qquad-E_{307}(\theta)+E_{308}(\theta)+E_{309}(\theta)+E_{310}(\theta)+E_{311}(\theta)+E_{312}(\theta)\right).$ (32)
$\theta$ $C_{3}(\theta)$ $\theta$ $C_{3}(\theta)$ $\theta$ $C_{3}(\theta)$
$0.5124$ $1.03763$ $0.5132$ $1.09167$ $0.5152$ $1.23885$
$0.5126$ $1.05089$ $0.5135$ $1.11171$ $0.5155$ $1.2675$
$0.5127$ $1.05778$ $0.514$ $1.14303$ $0.516$ $1.31988$
$0.5128$ $1.0652$ $0.5145$ $1.17429$ $0.5161$ $1.32981$
$0.513$ $1.07887$ $0.515$ $1.21983$ $\frac{16}{31}$ $1.3326$
6.4. Case 4. $\frac{16}{31}\leqslant\theta<\frac{14}{27}$
From here we note that Lemma 5.2 is not applicable. Buchstab’s identity yields
$S\left(\mathcal{A},(2x)^{\frac{1}{2}}\right)=$ $\ S\left(\mathcal{A},x^{\kappa}\right)-\sum_{\kappa\leqslant\alpha_{1}<\frac{1}{2}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)$
$=$ $\ S\left(\mathcal{A},x^{\kappa}\right)-\sum_{\kappa\leqslant\alpha_{1}<\tau}S\left(\mathcal{A}_{p_{1}},p_{1}\right)-\sum_{\tau\leqslant\alpha_{1}\leqslant\frac{3}{7}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)$
$-\sum_{\frac{3}{7}<\alpha_{1}<1-\theta}S\left(\mathcal{A}_{p_{1}},p_{1}\right)-\sum_{1-\theta\leqslant\alpha_{1}<\frac{1}{2}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)$
$=$ $\ S_{41}-S_{42}-S_{43}-S_{44}-S_{45}.$ (33)
Now, we can only use Lemma 5.3 to deal with parts of $S_{43}$ satisfy $\alpha_{1}\in\left[\frac{19\theta-7}{7},\frac{3}{7}\right]$, and discard the remaining parts. Combining with the loss integrals from $S_{42}$ and $S_{45}$, the total loss when $\frac{16}{31}\leqslant\theta<\frac{14}{27}$ is no more than
$\left(\int_{\kappa}^{\frac{\theta}{2}}\int_{1-\theta-t_{1}}^{\theta-t_{1}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{401}]\frac{\omega\left(\frac{t_{2}}{t_{1}}\right)\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}^{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{402}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in L_{403}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\int_{t_{4}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{404}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{405}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\tau}^{\frac{3}{7}}\texttt{Boole}[(t_{1})\in L_{406}]\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{3}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{4}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{409}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
$=$ $\ E_{401}(\theta)+E_{402}(\theta)-E_{403}(\theta)-E_{404}(\theta)+E_{405}(\theta)+E_{406}(\theta)+E_{407}(\theta)+E_{408}(\theta)+E_{409}(\theta),$ (34)
where
$L_{401}(\boldsymbol{\alpha}_{3}):=$ $\ \left\{\kappa\leqslant\alpha_{1}\leqslant\frac{\theta}{2},\ 1-\theta-\alpha_{1}\leqslant\alpha_{2}\leqslant\theta-\alpha_{1},\right$.
$\left.\quad\kappa\leqslant\alpha_{3}\leqslant\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\right\},$
$L_{402}(\boldsymbol{\alpha}_{3}):=$ $\ \left\{\kappa\leqslant\alpha_{1}<\tau,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{403}(\boldsymbol{\alpha}_{4}):=$ $\ \left\{\kappa\leqslant\alpha_{1}<\tau,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4},$
$\quad\alpha_{3}<\alpha_{4}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}),\ \boldsymbol{\alpha}_{4}\in\boldsymbol{G}_{4},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{404}(\boldsymbol{\alpha}_{5}):=$ $\ \left\{\kappa\leqslant\alpha_{1}<\tau,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4},$
$\quad\alpha_{3}<\alpha_{4}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}),\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},$
$\quad\alpha_{4}<\alpha_{5}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4}),\ \boldsymbol{\alpha}_{5}\in\boldsymbol{G}_{5},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{405}(\boldsymbol{\alpha}_{5}):=$ $\ \left\{\kappa\leqslant\alpha_{1}<\tau,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\boldsymbol{\alpha}_{3}\in\boldsymbol{U}_{3},\ (\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\in\boldsymbol{U}_{4},$
$\quad\kappa\leqslant\alpha_{4}<\min\left(\alpha_{3},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\right),\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},$
$\quad\kappa\leqslant\alpha_{5}<\min\left(\alpha_{4},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4})\right),\ \boldsymbol{\alpha}_{5}\notin\boldsymbol{G}_{5},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{406}(\alpha_{1}):=$ $\ \left\{\tau\leqslant\alpha_{1}\leqslant\frac{3}{7},\ \alpha_{1}\notin\left[\frac{19\theta-7}{7},\ \frac{3}{7}\right]\right\},$
$L_{409}(\boldsymbol{\alpha}_{3}):=$ $\ \left\{\tau\leqslant\alpha_{1}\leqslant\frac{3}{7},\ \alpha_{1}\in\left[\frac{19\theta-7}{7},\ \frac{3}{7}\right],\ \boldsymbol{\alpha}_{2}\notin\boldsymbol{G}_{2},\right$.
$\quad\kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\}.$
Finally, for $\frac{16}{31}\leqslant\theta<\frac{14}{27}$ we have
$C(\theta)\leqslant C_{4}(\theta):=1+\left(E_{401}(\theta)+E_{402}(\theta)-E_{403}(\theta)-E_{404}(\theta)+E_{405}(\theta)+E_{406}(\theta)+E_{407}(\theta)+E_{408}(\theta)+E_{409}(\theta)\right).$ (35)
$\theta$ $C_{4}(\theta)$ $\theta$ $C_{4}(\theta)$ $\theta$ $C_{4}(\theta)$
$0.5163$ $1.55978$ $0.5168$ $1.56977$ $0.5178$ $1.59006$
$0.5164$ $1.56181$ $0.5169$ $1.5718$ $0.518$ $1.59476$
$0.5165$ $1.56386$ $0.517$ $1.57396$ $0.5182$ $1.59927$
$0.5166$ $1.56604$ $0.5172$ $1.5775$ $0.5185$ $1.60315$
$0.5167$ $1.56778$ $0.5175$ $1.58362$ $\frac{14}{27}$ $1.60358$
6.5. Case 5. $\frac{14}{27}\leqslant\theta<\frac{11}{21}$
In this subrange, we have $\frac{50\theta-19}{17}\geqslant\frac{19\theta-7}{7}$, and the lower bound for $\alpha_{1}\in\boldsymbol{R}_{0}$ becomes $\frac{50\theta-19}{17}$. By an exactly same process as the decomposition in Case 4, the total loss when $\frac{14}{27}\leqslant\theta<\frac{11}{21}$ can be bounded by
$\left(\int_{\kappa}^{\frac{\theta}{2}}\int_{1-\theta-t_{1}}^{\theta-t_{1}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{501}]\frac{\omega\left(\frac{t_{2}}{t_{1}}\right)\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}^{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{502}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in L_{503}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\int_{t_{4}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{504}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{505}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\tau}^{\frac{3}{7}}\texttt{Boole}[(t_{1})\in L_{506}]\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{3}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{4}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{509}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
$=$ $\ E_{501}(\theta)+E_{502}(\theta)-E_{503}(\theta)-E_{504}(\theta)+E_{505}(\theta)+E_{506}(\theta)+E_{507}(\theta)+E_{508}(\theta)+E_{509}(\theta),$ (36)
where
$L_{501}(\boldsymbol{\alpha}_{3}):=$ $\ \left\{\kappa\leqslant\alpha_{1}\leqslant\frac{\theta}{2},\ 1-\theta-\alpha_{1}\leqslant\alpha_{2}\leqslant\theta-\alpha_{1},\right$.
$\left.\quad\kappa\leqslant\alpha_{3}\leqslant\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\right\},$
$L_{502}(\boldsymbol{\alpha}_{3}):=$ $\ \left\{\kappa\leqslant\alpha_{1}<\tau,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{503}(\boldsymbol{\alpha}_{4}):=$ $\ \left\{\kappa\leqslant\alpha_{1}<\tau,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4},$
$\quad\alpha_{3}<\alpha_{4}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}),\ \boldsymbol{\alpha}_{4}\in\boldsymbol{G}_{4},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{504}(\boldsymbol{\alpha}_{5}):=$ $\ \left\{\kappa\leqslant\alpha_{1}<\tau,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4},$
$\quad\alpha_{3}<\alpha_{4}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}),\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},$
$\quad\alpha_{4}<\alpha_{5}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4}),\ \boldsymbol{\alpha}_{5}\in\boldsymbol{G}_{5},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{505}(\boldsymbol{\alpha}_{5}):=$ $\ \left\{\kappa\leqslant\alpha_{1}<\tau,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\boldsymbol{\alpha}_{3}\in\boldsymbol{U}_{3},\ (\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\in\boldsymbol{U}_{4},$
$\quad\kappa\leqslant\alpha_{4}<\min\left(\alpha_{3},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\right),\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},$
$\quad\kappa\leqslant\alpha_{5}<\min\left(\alpha_{4},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4})\right),\ \boldsymbol{\alpha}_{5}\notin\boldsymbol{G}_{5},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{506}(\alpha_{1}):=$ $\ \left\{\tau\leqslant\alpha_{1}\leqslant\frac{3}{7},\ \alpha_{1}\notin\left[\frac{50\theta-19}{17},\ \frac{3}{7}\right]\right\},$
$L_{509}(\boldsymbol{\alpha}_{3}):=$ $\ \left\{\tau\leqslant\alpha_{1}\leqslant\frac{3}{7},\ \alpha_{1}\in\left[\frac{50\theta-19}{17},\ \frac{3}{7}\right],\ \boldsymbol{\alpha}_{2}\notin\boldsymbol{G}_{2},\right$.
$\quad\kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\}.$
Finally, for $\frac{14}{27}\leqslant\theta<\frac{11}{21}$ we have
$C(\theta)\leqslant C_{5}(\theta):=1+\left(E_{501}(\theta)+E_{502}(\theta)-E_{503}(\theta)-E_{504}(\theta)+E_{505}(\theta)+E_{506}(\theta)+E_{507}(\theta)+E_{508}(\theta)+E_{509}(\theta)\right).$ (37)
We remark that Harman [[16], Page 188] mentioned that $C(0.52)<1.632$, and we decide to use this value since our $C_{5}(0.52)$ is larger than $1.632$.
$\theta$ $C_{5}(\theta)$ $\theta$ $C_{5}(\theta)$ $\theta$ $C_{5}(\theta)$ $\theta$ $C_{5}(\theta)$
$0.5186$ $1.60737$ $0.52$ $1.632$ $0.5215$ $1.67861$ $0.5232$ $1.71925$
$0.5188$ $1.60886$ $0.5205$ $1.64649$ $0.522$ $1.68723$ $0.5234$ $1.72482$
$0.519$ $1.61355$ $0.5208$ $1.65321$ $0.5225$ $1.70046$ $0.5236$ $1.7305$
$0.5195$ $1.62485$ $0.5211$ $1.66246$ $0.523$ $1.71379$ $0.5237$ $1.7337$
6.6. Case 6. $\frac{11}{21}\leqslant\theta<\frac{92}{175}$
From here, Lemma 4.4 is not applicable. However, we do not need to worry about that. We note that for $\theta\geqslant\frac{11}{21}$, we must have $\boldsymbol{\alpha}_{2}\in A$ if $\alpha_{2}<\alpha_{1}<\tau=\frac{2}{7}$. If $\theta\leqslant\alpha_{1}+\alpha_{2}\leqslant 2\tau=\frac{4}{7}$, we have $\frac{3}{7}\leqslant 1-\alpha_{1}-\alpha_{2}<1-\theta$ and $\boldsymbol{\alpha}_{2}\in\boldsymbol{G}_{2}$. Otherwise we have $\alpha_{1}+\alpha_{2}<\theta$ ($\alpha_{2}<\alpha_{1}<\frac{\theta}{2}$), which means that $(\alpha_{1},\alpha_{2},2\theta-1)\in\boldsymbol{S}_{3}$ (one can easily check the corresponding conditions) and $\boldsymbol{\alpha}_{2}\in\boldsymbol{U}_{2}^{\prime}$ (or $\boldsymbol{U}_{2}$ when $\theta<\frac{7}{13}$). Thus, we can use Lemma 4.1 or Lemma 4.2 to deal with the sum
$\sum_{\kappa\leqslant\alpha_{2}<\alpha_{1}<\tau}S\left(\mathcal{A}_{p_{1}p_{2}},x^{\kappa}\right).$ (38)
The decomposing process of the remaining sums are similar to those in Case 5.
For the sake of simplicity, we choose to use the simple bound
$C(\theta)\leqslant 1.725+13.125(\theta-0.523)$ (39)
for $\frac{11}{21}\leqslant\theta<\frac{92}{175}$ given by Harman [[16], Page 186]. Note that $C(\theta)$ is monotonic increasing, we have
$C(\theta)\leqslant C\left(\frac{11}{21}\right)\leqslant 1.725+13.125\left(\frac{11}{21}-0.523\right)$ (40)
for $0.5237\leqslant\theta<\frac{11}{21}$. We also use the bound $C(0.525)<1.75$ (mentioned in [[16], Page 188]), which is slightly better than using (39) directly.
6.7. Case 7. $\frac{92}{175}\leqslant\theta<\frac{17}{32}$
From here, Lemma 5.3 is not applicable and we must discard the whole of
$\sum_{\tau\leqslant\alpha_{1}<\frac{3}{7}}S\left(\mathcal{A}_{p_{1}},p_{1}\right).$ (41)
For the sake of simplicity, we shall also use the simple bound
$C(\theta)\leqslant 1.725+13.125(\theta-0.523)$ (42)
for $\frac{92}{175}\leqslant\theta<\frac{17}{32}$ given by Harman [[16], Page 186].
6.8. Case 8. $\frac{17}{32}\leqslant\theta<\frac{7}{13}$
From here, Lemma 4.3 and Lemma 5.1 are not applicable. Buchstab’s identity yields
$S\left(\mathcal{A},(2x)^{\frac{1}{2}}\right)=$ $\ S\left(\mathcal{A},x^{\kappa}\right)-\sum_{\kappa\leqslant\alpha_{1}<\frac{1}{2}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)$
$=$ $\ S\left(\mathcal{A},x^{\kappa}\right)-\sum_{\kappa\leqslant\alpha_{1}<\tau}S\left(\mathcal{A}_{p_{1}},p_{1}\right)-\sum_{\tau\leqslant\alpha_{1}\leqslant\frac{3}{7}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)$
$-\sum_{\frac{3}{7}<\alpha_{1}<1-\theta}S\left(\mathcal{A}_{p_{1}},p_{1}\right)-\sum_{1-\theta\leqslant\alpha_{1}<\frac{1}{2}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)$
$=$ $\ S_{81}-S_{82}-S_{83}-S_{84}-S_{85}.$ (43)
$S_{81}$ and $S_{84}$ have asymptotic formulas. Since we cannot estimate the corresponding two-dimensional sums, we must discard the whole of $S_{83}$ and $S_{85}$. We can perform Buchstab’s identity twice more on $S_{82}$ to get
$S_{82}=$ $\ \sum_{\kappa\leqslant\alpha_{1}<\tau}S\left(\mathcal{A}_{p_{1}},p_{1}\right)$
$=$ $\ \sum_{\kappa\leqslant\alpha_{1}<\tau}S\left(\mathcal{A}_{p_{1}},x^{\kappa}\right)-\sum_{\begin{subarray}{c}\kappa\leqslant\alpha_{2}<\alpha_{1}<\tau\\ \boldsymbol{\alpha}_{2}\in\boldsymbol{G}_{2}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}},p_{2}\right)-\sum_{\begin{subarray}{c}\kappa\leqslant\alpha_{2}<\alpha_{1}<\tau\\ \boldsymbol{\alpha}_{2}\notin\boldsymbol{G}_{2}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}},x^{\kappa}\right)$
$+\sum_{\begin{subarray}{c}\kappa\leqslant\alpha_{2}<\alpha_{1}<\tau\\ \boldsymbol{\alpha}_{2}\notin\boldsymbol{G}_{2}\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\in\boldsymbol{G}_{3}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}},p_{3}\right)+\sum_{\begin{subarray}{c}\kappa\leqslant\alpha_{2}<\alpha_{1}<\tau\\ \boldsymbol{\alpha}_{2}\notin\boldsymbol{G}_{2}\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}},p_{3}\right)$
$=$ $\ S_{821}-S_{822}-S_{823}+S_{824}+S_{825}.$ (44)
$S_{822}$ and $S_{824}$ have asymptotic formulas. We can use Lemma 4.2 to give an asymptotic formula for $S_{823}$ (and also $S_{821}$) by the discussion in Case 6. For $S_{825}$ we can check the same conditions as those in Case 1: if $\boldsymbol{\alpha}_{3}\in\boldsymbol{U}_{3}$ and
$(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\in\boldsymbol{U}_{4},$
for parts of $S_{825}$, then we can use Buchstab’s identity twice again to get a five-dimensional sum. We can even get a seven-dimensional sum if we have $\boldsymbol{\alpha}_{5}\in\boldsymbol{U}_{5}$ and
$(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5},\alpha_{5})\in\boldsymbol{U}_{6}.$
For the sums without further decompositions, we use the reversed Buchstab’s identity to gain savings from almost-primes counted. Note that the halved value of $\kappa$ in this subrange of $\theta$ doubles the maximum number of prime factors of the elements counted in these sums. Summing up those resulted integrals, the total loss when $\frac{17}{32}\leqslant\theta<\frac{7}{13}$ is no more than
$\left(\int_{1-\theta}^{\frac{1}{2}}\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{802}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in L_{803}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\int_{t_{4}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{804}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{805}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\int_{t_{5}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6})\in L_{806}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{t_{6}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}^{2}}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\tau}^{\frac{3}{7}}\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
$=$ $\ E_{801}(\theta)+E_{802}(\theta)-E_{803}(\theta)-E_{804}(\theta)+E_{805}(\theta)-E_{806}(\theta)+E_{807}(\theta),$ (45)
where
$L_{802}(\boldsymbol{\alpha}_{3}):=$ $\ \left\{\kappa\leqslant\alpha_{1}<\tau,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{803}(\boldsymbol{\alpha}_{4}):=$ $\ \left\{\kappa\leqslant\alpha_{1}<\tau,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4},$
$\quad\alpha_{3}<\alpha_{4}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}),\ \boldsymbol{\alpha}_{4}\in\boldsymbol{G}_{4},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{804}(\boldsymbol{\alpha}_{5}):=$ $\ \left\{\kappa\leqslant\alpha_{1}<\tau,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4},$
$\quad\alpha_{3}<\alpha_{4}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}),\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},$
$\quad\alpha_{4}<\alpha_{5}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4}),\ \boldsymbol{\alpha}_{5}\in\boldsymbol{G}_{5},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{805}(\boldsymbol{\alpha}_{5}):=$ $\ \left\{\kappa\leqslant\alpha_{1}<\tau,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\boldsymbol{\alpha}_{3}\in\boldsymbol{U}_{3},\ (\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\in\boldsymbol{U}_{4},$
$\quad\kappa\leqslant\alpha_{4}<\min\left(\alpha_{3},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\right),\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},$
$\quad\kappa\leqslant\alpha_{5}<\min\left(\alpha_{4},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4})\right),\ \boldsymbol{\alpha}_{5}\notin\boldsymbol{G}_{5},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{806}(\boldsymbol{\alpha}_{6}):=$ $\ \left\{\kappa\leqslant\alpha_{1}<\tau,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\boldsymbol{\alpha}_{3}\in\boldsymbol{U}_{3},\ (\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\in\boldsymbol{U}_{4},$
$\quad\kappa\leqslant\alpha_{4}<\min\left(\alpha_{3},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\right),\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},$
$\quad\kappa\leqslant\alpha_{5}<\min\left(\alpha_{4},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4})\right),\ \boldsymbol{\alpha}_{5}\notin\boldsymbol{G}_{5},$
$\quad\alpha_{5}<\alpha_{6}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4}-\alpha_{5}),\ \boldsymbol{\alpha}_{6}\in\boldsymbol{G}_{6},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\}.$
Finally, for $\frac{17}{32}\leqslant\theta<\frac{7}{13}$ we have
$C(\theta)\leqslant C_{8}(\theta):=1+\left(E_{801}(\theta)+E_{802}(\theta)-E_{803}(\theta)-E_{804}(\theta)+E_{805}(\theta)-E_{806}(\theta)+E_{807}(\theta)\right).$ (46)
$\theta$ $C_{8}(\theta)$ $\theta$ $C_{8}(\theta)$ $\theta$ $C_{8}(\theta)$ $\theta$ $C_{8}(\theta)$
$0.5316$ $1.8603$ $0.5342$ $1.87985$ $0.5353$ $1.90671$ $0.5373$ $1.93236$
$0.5328$ $1.86462$ $0.5343$ $1.89006$ $0.5355$ $1.91223$ $0.5374$ $1.93399$
$0.533$ $1.86571$ $0.5345$ $1.89323$ $0.5358$ $1.91673$ $0.5376$ $1.93644$
$0.5333$ $1.86834$ $0.5348$ $1.8991$ $0.536$ $1.91998$ $0.5377$ $1.93829$
$0.5335$ $1.87103$ $0.5349$ $1.90152$ $0.5363$ $1.92343$ $0.5378$ $1.94163$
$0.534$ $1.87798$ $0.5351$ $1.90373$ $0.5365$ $1.92673$ $0.5381$ $1.94862$
$0.5341$ $1.87885$ $0.5352$ $1.90524$ $0.537$ $1.92762$ $0.5384$ $1.94911$
6.9. Case 9. $\frac{7}{13}\leqslant\theta<\frac{6}{11}$
From here, ${U}_{j}$ and ${U}_{j}^{\prime}$ do not cover the same region. Since we have $\tau^{\prime}=\frac{5-6\theta}{7}<\frac{2}{7}=\tau$ and $\kappa^{\prime}=\frac{11-20\theta}{6}<\frac{3-5\theta}{7}=\kappa$ in this subrange, Buchstab’s identity gives us
$S\left(\mathcal{A},(2x)^{\frac{1}{2}}\right)=$ $\ S\left(\mathcal{A},x^{\kappa}\right)-\sum_{\kappa\leqslant\alpha_{1}<\frac{1}{2}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)$
$=$ $\ S\left(\mathcal{A},x^{\kappa}\right)-\sum_{\kappa\leqslant\alpha_{1}<\tau^{\prime}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)-\sum_{\tau^{\prime}\leqslant\alpha_{1}<\tau}S\left(\mathcal{A}_{p_{1}},p_{1}\right)-\sum_{\tau\leqslant\alpha_{1}\leqslant\frac{3}{7}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)$
$-\sum_{\frac{3}{7}<\alpha_{1}<1-\theta}S\left(\mathcal{A}_{p_{1}},p_{1}\right)-\sum_{1-\theta\leqslant\alpha_{1}<\frac{1}{2}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)$
$=$ $\ S_{91}-S_{92}-S_{93}-S_{94}-S_{95}-S_{96}.$ (47)
We give asymptotic formulas for $S_{91}$ and $S_{95}$, and discard $S_{94}$ and $S_{96}$. For $S_{92}$ and $S_{93}$, we need to check several conditions. For $S_{92}$ we have $\alpha_{2}<\alpha_{1}<\frac{5-6\theta}{7}$. Since $\frac{8\theta-2}{7}>\frac{5-6\theta}{7}$, we have $\boldsymbol{\alpha}_{2}\in\boldsymbol{T}_{2}^{*}$ and $\boldsymbol{\alpha}_{2}\in\boldsymbol{U}_{2}$. Now we can use Buchstab’s identity to get
$S_{92}=$ $\ \sum_{\kappa\leqslant\alpha_{1}<\tau^{\prime}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)$
$=$ $\ \sum_{\kappa\leqslant\alpha_{1}<\tau^{\prime}}S\left(\mathcal{A}_{p_{1}},x^{\kappa}\right)-\sum_{\begin{subarray}{c}\kappa\leqslant\alpha_{2}<\alpha_{1}<\tau^{\prime}\\ \boldsymbol{\alpha}_{2}\in\boldsymbol{G}_{2}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}},p_{2}\right)-\sum_{\begin{subarray}{c}\kappa\leqslant\alpha_{2}<\alpha_{1}<\tau^{\prime}\\ \boldsymbol{\alpha}_{2}\notin\boldsymbol{G}_{2}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}},x^{\kappa}\right)$
$+\sum_{\begin{subarray}{c}\kappa\leqslant\alpha_{2}<\alpha_{1}<\tau^{\prime}\\ \boldsymbol{\alpha}_{2}\notin\boldsymbol{G}_{2}\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\in\boldsymbol{G}_{3}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}},p_{3}\right)+\sum_{\begin{subarray}{c}\kappa\leqslant\alpha_{2}<\alpha_{1}<\tau^{\prime}\\ \boldsymbol{\alpha}_{2}\notin\boldsymbol{G}_{2}\\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}},p_{3}\right)$
$=$ $\ S_{921}-S_{922}-S_{923}+S_{924}+S_{925}.$ (48)
Now we can find the loss from $S_{92}$ by the exactly same process used in $S_{82}$. For $S_{93}$ we shall use a similar process. Since $\alpha_{1}\geqslant\tau^{\prime}$ in $S_{93}$, we cannot ensure that $\boldsymbol{\alpha}_{2}\in\boldsymbol{T}_{2}^{*}$ after a Buchstab iteration. However, by the discussion in Case 6, we only need to consider the case $\alpha_{2}<\alpha_{1}<\frac{\theta}{2}$. Since $\theta<\frac{6}{11}$, we find that $\left(\frac{\theta}{2},\frac{\theta}{2},2\theta-1\right)\in\boldsymbol{S}_{3}$, hence $(\alpha_{1},\alpha_{2},2\theta-1)\in\boldsymbol{S}_{3}$ and $\boldsymbol{\alpha}_{2}\in\boldsymbol{U}_{2}^{\prime}$. By a similar Buchstab iteration process (with $\kappa^{\prime}$ and $\tau^{\prime}$), we get
$S_{93}=$ $\ \sum_{\tau^{\prime}\leqslant\alpha_{1}<\tau}S\left(\mathcal{A}_{p_{1}},p_{1}\right)$
$=$ $\ \sum_{\tau^{\prime}\leqslant\alpha_{1}<\tau}S\left(\mathcal{A}_{p_{1}},x^{\kappa^{\prime}}\right)-\sum_{\begin{subarray}{c}\tau^{\prime}\leqslant\alpha_{1}<\tau\\ \kappa^{\prime}\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\\ \boldsymbol{\alpha}_{2}\in\boldsymbol{G}_{2}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}},p_{2}\right)-\sum_{\begin{subarray}{c}\tau^{\prime}\leqslant\alpha_{1}<\tau\\ \kappa^{\prime}\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\\ \boldsymbol{\alpha}_{2}\notin\boldsymbol{G}_{2}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}},x^{\kappa^{\prime}}\right)$
$+\sum_{\begin{subarray}{c}\tau^{\prime}\leqslant\alpha_{1}<\tau\\ \kappa^{\prime}\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\\ \boldsymbol{\alpha}_{2}\notin\boldsymbol{G}_{2}\\ \kappa^{\prime}\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\in\boldsymbol{G}_{3}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}},p_{3}\right)+\sum_{\begin{subarray}{c}\tau^{\prime}\leqslant\alpha_{1}<\tau\\ \kappa^{\prime}\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\\ \boldsymbol{\alpha}_{2}\notin\boldsymbol{G}_{2}\\ \kappa^{\prime}\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right)\\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3}\end{subarray}}S\left(\mathcal{A}_{p_{1}p_{2}p_{3}},p_{3}\right)$
$=$ $\ S_{931}-S_{932}-S_{933}+S_{934}+S_{935}.$ (49)
We can find the loss from $S_{93}$ by almost the same process as used in $S_{92}$. We need to replace many $\kappa$ with $\kappa^{\prime}$ and $\boldsymbol{U}_{j}$ with $\boldsymbol{U}_{j}^{\prime}$ in the resulting integrals. We note that seven-dimensional sums are considered in both $S_{92}$ and $S_{93}$. Putting those integrals together, the total loss when $\frac{7}{13}\leqslant\theta<\frac{6}{11}$ can be bounded by
$\left(\int_{1-\theta}^{\frac{1}{2}}\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{902}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in L_{903}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\int_{t_{4}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{904}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{905}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\int_{t_{5}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6})\in L_{906}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{t_{6}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}^{2}}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\quad\quad\quad\int_{t_{5}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}}\int_{t_{6}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{2}}$
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{907}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\quad\quad\quad\int_{\kappa}^{\min\left(t_{5},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}\right)}\int_{\kappa}^{\min\left(t_{6},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{2}\right)}$
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{908}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{909}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in L_{910}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\int_{t_{4}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{911}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa^{\prime}}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa^{\prime}}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{912}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa^{\prime}}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa^{\prime}}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\int_{t_{5}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6})\in L_{913}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{t_{6}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}^{2}}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa^{\prime}}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa^{\prime}}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\quad\quad\quad\int_{t_{5}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}}\int_{t_{6}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{2}}$
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{914}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa^{\prime}}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa^{\prime}}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\quad\quad\quad\int_{\kappa^{\prime}}^{\min\left(t_{5},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}\right)}\int_{\kappa^{\prime}}^{\min\left(t_{6},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{2}\right)}$
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{915}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\tau}^{\frac{3}{7}}\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
$=$ $\ E_{901}(\theta)+E_{902}(\theta)-E_{903}(\theta)-E_{904}(\theta)+E_{905}(\theta)-E_{906}(\theta)-E_{907}(\theta)+E_{908}(\theta)$
$+E_{909}(\theta)-E_{910}(\theta)-E_{911}(\theta)+E_{912}(\theta)-E_{913}(\theta)-E_{914}(\theta)+E_{915}(\theta)+E_{916}(\theta),$ (50)
where
$L_{902}(\boldsymbol{\alpha}_{3}):=$ $\ \left\{\kappa\leqslant\alpha_{1}<\tau^{\prime},\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{903}(\boldsymbol{\alpha}_{4}):=$ $\ \left\{\kappa\leqslant\alpha_{1}<\tau^{\prime},\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4},$
$\quad\alpha_{3}<\alpha_{4}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}),\ \boldsymbol{\alpha}_{4}\in\boldsymbol{G}_{4},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{904}(\boldsymbol{\alpha}_{5}):=$ $\ \left\{\kappa\leqslant\alpha_{1}<\tau^{\prime},\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4},$
$\quad\alpha_{3}<\alpha_{4}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}),\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},$
$\quad\alpha_{4}<\alpha_{5}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4}),\ \boldsymbol{\alpha}_{5}\in\boldsymbol{G}_{5},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{905}(\boldsymbol{\alpha}_{5}):=$ $\ \left\{\kappa\leqslant\alpha_{1}<\tau^{\prime},\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\boldsymbol{\alpha}_{3}\in\boldsymbol{U}_{3},\ (\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\in\boldsymbol{U}_{4},$
$\quad\kappa\leqslant\alpha_{4}<\min\left(\alpha_{3},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\right),\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},$
$\quad\kappa\leqslant\alpha_{5}<\min\left(\alpha_{4},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4})\right),\ \boldsymbol{\alpha}_{5}\notin\boldsymbol{G}_{5},$
$\quad\text{either }\boldsymbol{\alpha}_{5}\notin\boldsymbol{U}_{5}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5},\alpha_{5})\notin\boldsymbol{U}_{6},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{906}(\boldsymbol{\alpha}_{6}):=$ $\ \left\{\kappa\leqslant\alpha_{1}<\tau^{\prime},\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\boldsymbol{\alpha}_{3}\in\boldsymbol{U}_{3},\ (\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\in\boldsymbol{U}_{4},$
$\quad\kappa\leqslant\alpha_{4}<\min\left(\alpha_{3},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\right),\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},$
$\quad\kappa\leqslant\alpha_{5}<\min\left(\alpha_{4},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4})\right),\ \boldsymbol{\alpha}_{5}\notin\boldsymbol{G}_{5},$
$\quad\text{either }\boldsymbol{\alpha}_{5}\notin\boldsymbol{U}_{5}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5},\alpha_{5})\notin\boldsymbol{U}_{6},$
$\quad\alpha_{5}<\alpha_{6}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4}-\alpha_{5}),\ \boldsymbol{\alpha}_{6}\in\boldsymbol{G}_{6},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{907}(\boldsymbol{\alpha}_{7}):=$ $\ \left\{\kappa\leqslant\alpha_{1}<\tau^{\prime},\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\boldsymbol{\alpha}_{3}\in\boldsymbol{U}_{3},\ (\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\in\boldsymbol{U}_{4},$
$\quad\kappa\leqslant\alpha_{4}<\min\left(\alpha_{3},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\right),\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},$
$\quad\kappa\leqslant\alpha_{5}<\min\left(\alpha_{4},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4})\right),\ \boldsymbol{\alpha}_{5}\notin\boldsymbol{G}_{5},$
$\quad\text{either }\boldsymbol{\alpha}_{5}\notin\boldsymbol{U}_{5}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5},\alpha_{5})\notin\boldsymbol{U}_{6},$
$\quad\alpha_{5}<\alpha_{6}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4}-\alpha_{5}),\ \boldsymbol{\alpha}_{6}\notin\boldsymbol{G}_{6},$
$\quad\alpha_{6}<\alpha_{7}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4}-\alpha_{5}-\alpha_{6}),\ \boldsymbol{\alpha}_{7}\in\boldsymbol{G}_{7},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{908}(\boldsymbol{\alpha}_{7}):=$ $\ \left\{\kappa\leqslant\alpha_{1}<\tau^{\prime},\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\boldsymbol{\alpha}_{3}\in\boldsymbol{U}_{3},\ (\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\in\boldsymbol{U}_{4},$
$\quad\kappa\leqslant\alpha_{4}<\min\left(\alpha_{3},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\right),\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},$
$\quad\kappa\leqslant\alpha_{5}<\min\left(\alpha_{4},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4})\right),\ \boldsymbol{\alpha}_{5}\notin\boldsymbol{G}_{5},$
$\quad\boldsymbol{\alpha}_{5}\in\boldsymbol{U}_{5},\ (\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5},\alpha_{5})\in\boldsymbol{U}_{6},$
$\quad\kappa\leqslant\alpha_{6}<\min\left(\alpha_{5},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4}-\alpha_{5})\right),\ \boldsymbol{\alpha}_{6}\notin\boldsymbol{G}_{6},$
$\quad\kappa\leqslant\alpha_{7}<\min\left(\alpha_{6},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4}-\alpha_{5}-\alpha_{6})\right),\ \boldsymbol{\alpha}_{7}\notin\boldsymbol{G}_{7},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{909}(\boldsymbol{\alpha}_{3}):=$ $\ \left\{\tau^{\prime}\leqslant\alpha_{1}<\tau,\ \kappa^{\prime}\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}^{\prime}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}^{\prime}_{4},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa^{\prime}\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{910}(\boldsymbol{\alpha}_{4}):=$ $\ \left\{\tau^{\prime}\leqslant\alpha_{1}<\tau,\ \kappa^{\prime}\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}^{\prime}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}^{\prime}_{4},$
$\quad\alpha_{3}<\alpha_{4}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}),\ \boldsymbol{\alpha}_{4}\in\boldsymbol{G}_{4},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa^{\prime}\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{911}(\boldsymbol{\alpha}_{5}):=$ $\ \left\{\tau^{\prime}\leqslant\alpha_{1}<\tau,\ \kappa^{\prime}\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}^{\prime}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}^{\prime}_{4},$
$\quad\alpha_{3}<\alpha_{4}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}),\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},$
$\quad\alpha_{4}<\alpha_{5}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4}),\ \boldsymbol{\alpha}_{5}\in\boldsymbol{G}_{5},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa^{\prime}\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{912}(\boldsymbol{\alpha}_{5}):=$ $\ \left\{\tau^{\prime}\leqslant\alpha_{1}<\tau,\ \kappa^{\prime}\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\boldsymbol{\alpha}_{3}\in\boldsymbol{U}^{\prime}_{3},\ (\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\in\boldsymbol{U}^{\prime}_{4},$
$\quad\kappa^{\prime}\leqslant\alpha_{4}<\min\left(\alpha_{3},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\right),\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},$
$\quad\kappa^{\prime}\leqslant\alpha_{5}<\min\left(\alpha_{4},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4})\right),\ \boldsymbol{\alpha}_{5}\notin\boldsymbol{G}_{5},$
$\quad\text{either }\boldsymbol{\alpha}_{5}\notin\boldsymbol{U}^{\prime}_{5}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5},\alpha_{5})\notin\boldsymbol{U}^{\prime}_{6},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa^{\prime}\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{913}(\boldsymbol{\alpha}_{6}):=$ $\ \left\{\tau^{\prime}\leqslant\alpha_{1}<\tau,\ \kappa^{\prime}\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\boldsymbol{\alpha}_{3}\in\boldsymbol{U}^{\prime}_{3},\ (\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\in\boldsymbol{U}^{\prime}_{4},$
$\quad\kappa^{\prime}\leqslant\alpha_{4}<\min\left(\alpha_{3},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\right),\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},$
$\quad\kappa^{\prime}\leqslant\alpha_{5}<\min\left(\alpha_{4},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4})\right),\ \boldsymbol{\alpha}_{5}\notin\boldsymbol{G}_{5},$
$\quad\text{either }\boldsymbol{\alpha}_{5}\notin\boldsymbol{U}^{\prime}_{5}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5},\alpha_{5})\notin\boldsymbol{U}^{\prime}_{6},$
$\quad\alpha_{5}<\alpha_{6}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4}-\alpha_{5}),\ \boldsymbol{\alpha}_{6}\in\boldsymbol{G}_{6},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa^{\prime}\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{914}(\boldsymbol{\alpha}_{7}):=$ $\ \left\{\tau^{\prime}\leqslant\alpha_{1}<\tau,\ \kappa^{\prime}\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\boldsymbol{\alpha}_{3}\in\boldsymbol{U}^{\prime}_{3},\ (\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\in\boldsymbol{U}^{\prime}_{4},$
$\quad\kappa^{\prime}\leqslant\alpha_{4}<\min\left(\alpha_{3},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\right),\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},$
$\quad\kappa^{\prime}\leqslant\alpha_{5}<\min\left(\alpha_{4},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4})\right),\ \boldsymbol{\alpha}_{5}\notin\boldsymbol{G}_{5},$
$\quad\text{either }\boldsymbol{\alpha}_{5}\notin\boldsymbol{U}^{\prime}_{5}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5},\alpha_{5})\notin\boldsymbol{U}^{\prime}_{6},$
$\quad\alpha_{5}<\alpha_{6}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4}-\alpha_{5}),\ \boldsymbol{\alpha}_{6}\notin\boldsymbol{G}_{6},$
$\quad\alpha_{6}<\alpha_{7}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4}-\alpha_{5}-\alpha_{6}),\ \boldsymbol{\alpha}_{7}\in\boldsymbol{G}_{7},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa^{\prime}\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{915}(\boldsymbol{\alpha}_{7}):=$ $\ \left\{\tau^{\prime}\leqslant\alpha_{1}<\tau,\ \kappa^{\prime}\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\boldsymbol{\alpha}_{3}\in\boldsymbol{U}^{\prime}_{3},\ (\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\in\boldsymbol{U}^{\prime}_{4},$
$\quad\kappa^{\prime}\leqslant\alpha_{4}<\min\left(\alpha_{3},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\right),\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},$
$\quad\kappa^{\prime}\leqslant\alpha_{5}<\min\left(\alpha_{4},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4})\right),\ \boldsymbol{\alpha}_{5}\notin\boldsymbol{G}_{5},$
$\quad\boldsymbol{\alpha}_{5}\in\boldsymbol{U}^{\prime}_{5},\ (\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5},\alpha_{5})\in\boldsymbol{U}^{\prime}_{6},$
$\quad\kappa^{\prime}\leqslant\alpha_{6}<\min\left(\alpha_{5},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4}-\alpha_{5})\right),\ \boldsymbol{\alpha}_{6}\notin\boldsymbol{G}_{6},$
$\quad\kappa^{\prime}\leqslant\alpha_{7}<\min\left(\alpha_{6},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4}-\alpha_{5}-\alpha_{6})\right),\ \boldsymbol{\alpha}_{7}\notin\boldsymbol{G}_{7},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa^{\prime}\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\}.$
Finally, for $\frac{7}{13}\leqslant\theta<\frac{6}{11}$ we have
$C(\theta)\leqslant C_{9}(\theta):=$ $\ 1+\left(E_{901}(\theta)+E_{902}(\theta)-E_{903}(\theta)-E_{904}(\theta)+E_{905}(\theta)-E_{906}(\theta)-E_{907}(\theta)+E_{908}(\theta)\right$.
$\left.\qquad+E_{909}(\theta)-E_{910}(\theta)-E_{911}(\theta)+E_{912}(\theta)-E_{913}(\theta)-E_{914}(\theta)+E_{915}(\theta)+E_{916}(\theta)\right).$ (51)
$\theta$ $C_{9}(\theta)$ $\theta$ $C_{9}(\theta)$ $\theta$ $C_{9}(\theta)$ $\theta$ $C_{9}(\theta)$
$0.5385$ $1.95009$ $0.5402$ $2.00356$ $0.5422$ $2.08389$ $0.544$ $2.16035$
$0.5388$ $1.95279$ $0.5405$ $2.023$ $0.5425$ $2.10289$ $0.5442$ $2.16134$
$0.539$ $1.95677$ $0.5408$ $2.03985$ $0.5428$ $2.10903$ $0.5445$ $2.18608$
$0.5392$ $1.95916$ $0.541$ $2.05815$ $0.543$ $2.12223$ $0.5448$ $2.20478$
$0.5395$ $1.96639$ $0.5412$ $2.06933$ $0.5432$ $2.12238$ $0.545$ $2.2179$
$0.5398$ $1.986$ $0.5415$ $2.0785$ $0.5435$ $2.13851$ $0.5452$ $2.22554$
$0.54$ $1.99921$ $0.542$ $2.07929$ $0.5438$ $2.15308$ $0.5454$ $2.22589$
6.10. Case 10. $\frac{6}{11}\leqslant\theta\leqslant 0.679$
In the last case, we cannot ensure that $(\alpha_{1},\alpha_{2},2\theta-1)\in\boldsymbol{S}_{3}$ since $\theta\geqslant\frac{6}{11}$. Using Buchstab’s identity, we have
$S\left(\mathcal{A},(2x)^{\frac{1}{2}}\right)=$ $\ S\left(\mathcal{A},x^{\kappa}\right)-\sum_{\kappa\leqslant\alpha_{1}<\frac{1}{2}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)$
$=$ $\ S\left(\mathcal{A},x^{\kappa}\right)-\sum_{\kappa\leqslant\alpha_{1}<\tau}S\left(\mathcal{A}_{p_{1}},p_{1}\right)-\sum_{\tau\leqslant\alpha_{1}\leqslant\frac{3}{7}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)$
$-\sum_{\frac{3}{7}<\alpha_{1}<1-\theta}S\left(\mathcal{A}_{p_{1}},p_{1}\right)-\sum_{1-\theta\leqslant\alpha_{1}<\frac{1}{2}}S\left(\mathcal{A}_{p_{1}},p_{1}\right)$
$=$ $\ S_{X1}-S_{X2}-S_{X3}-S_{X4}-S_{X5}.$ (52)
The next decomposing process is quite similar to the process in Case 9. The only difference is that we do not need to consider
$E_{909}(\theta)-E_{910}(\theta)-E_{911}(\theta)+E_{912}(\theta)-E_{913}(\theta)-E_{914}(\theta)+E_{915}(\theta)$
since $\tau=\tau^{\prime}=\frac{5-6\theta}{7}$ now. The total loss when $\frac{6}{11}\leqslant\theta<\frac{4}{7}$ is no more than
$\left(\int_{1-\theta}^{\frac{1}{2}}\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{X02}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in L_{X03}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\int_{t_{4}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{X04}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{X05}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\int_{t_{5}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6})\in L_{X06}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{t_{6}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}^{2}}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\quad\quad\quad\int_{t_{5}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}}\int_{t_{6}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{2}}$
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{X07}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\quad\quad\quad\int_{\kappa}^{\min\left(t_{5},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}\right)}\int_{\kappa}^{\min\left(t_{6},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{2}\right)}$
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{X08}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\tau}^{\frac{3}{7}}\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
$=$ $\ E_{X01}(\theta)+E_{X02}(\theta)-E_{X03}(\theta)-E_{X04}(\theta)+E_{X05}(\theta)-E_{X06}(\theta)-E_{X07}(\theta)+E_{X08}(\theta)+E_{X09}(\theta),$ (53)
where
$L_{X02}(\boldsymbol{\alpha}_{3}):=$ $\ \left\{\kappa\leqslant\alpha_{1}<\tau,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{X03}(\boldsymbol{\alpha}_{4}):=$ $\ \left\{\kappa\leqslant\alpha_{1}<\tau,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4},$
$\quad\alpha_{3}<\alpha_{4}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}),\ \boldsymbol{\alpha}_{4}\in\boldsymbol{G}_{4},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{X04}(\boldsymbol{\alpha}_{5}):=$ $\ \left\{\kappa\leqslant\alpha_{1}<\tau,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\text{either }\boldsymbol{\alpha}_{3}\notin\boldsymbol{U}_{3}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\notin\boldsymbol{U}_{4},$
$\quad\alpha_{3}<\alpha_{4}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}),\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},$
$\quad\alpha_{4}<\alpha_{5}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4}),\ \boldsymbol{\alpha}_{5}\in\boldsymbol{G}_{5},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{X05}(\boldsymbol{\alpha}_{5}):=$ $\ \left\{\kappa\leqslant\alpha_{1}<\tau,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\boldsymbol{\alpha}_{3}\in\boldsymbol{U}_{3},\ (\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\in\boldsymbol{U}_{4},$
$\quad\kappa\leqslant\alpha_{4}<\min\left(\alpha_{3},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\right),\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},$
$\quad\kappa\leqslant\alpha_{5}<\min\left(\alpha_{4},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4})\right),\ \boldsymbol{\alpha}_{5}\notin\boldsymbol{G}_{5},$
$\quad\text{either }\boldsymbol{\alpha}_{5}\notin\boldsymbol{U}_{5}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5},\alpha_{5})\notin\boldsymbol{U}_{6},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{X06}(\boldsymbol{\alpha}_{6}):=$ $\ \left\{\kappa\leqslant\alpha_{1}<\tau,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\boldsymbol{\alpha}_{3}\in\boldsymbol{U}_{3},\ (\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\in\boldsymbol{U}_{4},$
$\quad\kappa\leqslant\alpha_{4}<\min\left(\alpha_{3},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\right),\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},$
$\quad\kappa\leqslant\alpha_{5}<\min\left(\alpha_{4},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4})\right),\ \boldsymbol{\alpha}_{5}\notin\boldsymbol{G}_{5},$
$\quad\text{either }\boldsymbol{\alpha}_{5}\notin\boldsymbol{U}_{5}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5},\alpha_{5})\notin\boldsymbol{U}_{6},$
$\quad\alpha_{5}<\alpha_{6}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4}-\alpha_{5}),\ \boldsymbol{\alpha}_{6}\in\boldsymbol{G}_{6},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{X07}(\boldsymbol{\alpha}_{7}):=$ $\ \left\{\kappa\leqslant\alpha_{1}<\tau,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\boldsymbol{\alpha}_{3}\in\boldsymbol{U}_{3},\ (\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\in\boldsymbol{U}_{4},$
$\quad\kappa\leqslant\alpha_{4}<\min\left(\alpha_{3},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\right),\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},$
$\quad\kappa\leqslant\alpha_{5}<\min\left(\alpha_{4},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4})\right),\ \boldsymbol{\alpha}_{5}\notin\boldsymbol{G}_{5},$
$\quad\text{either }\boldsymbol{\alpha}_{5}\notin\boldsymbol{U}_{5}\text{ or }(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5},\alpha_{5})\notin\boldsymbol{U}_{6},$
$\quad\alpha_{5}<\alpha_{6}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4}-\alpha_{5}),\ \boldsymbol{\alpha}_{6}\notin\boldsymbol{G}_{6},$
$\quad\alpha_{6}<\alpha_{7}<\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4}-\alpha_{5}-\alpha_{6}),\ \boldsymbol{\alpha}_{7}\in\boldsymbol{G}_{7},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\},$
$L_{X08}(\boldsymbol{\alpha}_{7}):=$ $\ \left\{\kappa\leqslant\alpha_{1}<\tau,\ \kappa\leqslant\alpha_{3}<\min\left(\alpha_{2},\frac{1}{2}(1-\alpha_{1}-\alpha_{2})\right),\ \boldsymbol{\alpha}_{3}\notin\boldsymbol{G}_{3},\right$.
$\quad\boldsymbol{\alpha}_{3}\in\boldsymbol{U}_{3},\ (\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{3})\in\boldsymbol{U}_{4},$
$\quad\kappa\leqslant\alpha_{4}<\min\left(\alpha_{3},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3})\right),\ \boldsymbol{\alpha}_{4}\notin\boldsymbol{G}_{4},$
$\quad\kappa\leqslant\alpha_{5}<\min\left(\alpha_{4},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4})\right),\ \boldsymbol{\alpha}_{5}\notin\boldsymbol{G}_{5},$
$\quad\boldsymbol{\alpha}_{5}\in\boldsymbol{U}_{5},\ (\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5},\alpha_{5})\in\boldsymbol{U}_{6},$
$\quad\kappa\leqslant\alpha_{6}<\min\left(\alpha_{5},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4}-\alpha_{5})\right),\ \boldsymbol{\alpha}_{6}\notin\boldsymbol{G}_{6},$
$\quad\kappa\leqslant\alpha_{7}<\min\left(\alpha_{6},\frac{1}{2}(1-\alpha_{1}-\alpha_{2}-\alpha_{3}-\alpha_{4}-\alpha_{5}-\alpha_{6})\right),\ \boldsymbol{\alpha}_{7}\notin\boldsymbol{G}_{7},$
$\left.\quad\kappa\leqslant\alpha_{1}<\frac{1}{2},\ \kappa\leqslant\alpha_{2}<\min\left(\alpha_{1},\frac{1}{2}(1-\alpha_{1})\right)\right\}.$
For $\frac{6}{11}\leqslant\theta<\frac{4}{7}$ we have
$C(\theta)\leqslant C_{X}(\theta):=$ $\ 1+\left(E_{X01}(\theta)+E_{X02}(\theta)-E_{X03}(\theta)-E_{X04}(\theta)\right$.
$\left.\qquad+E_{X05}(\theta)-E_{X06}(\theta)-E_{X07}(\theta)+E_{X08}(\theta)+E_{X09}(\theta)\right).$ (54)
$\theta$ $C_{X}(\theta)$ $\theta$ $C_{X}(\theta)$ $\theta$ $C_{X}(\theta)$
$0.546$ $2.33807$ $0.553$ $2.44031$ $0.561$ $2.68948$
$0.547$ $2.35981$ $0.555$ $2.45396$ $0.562$ $2.71434$
$0.549$ $2.3809$ $0.556$ $2.50757$ $0.563$ $2.80532$
$0.55$ $2.39955$ $0.558$ $2.51193$ $0.564$ $2.87871$
$0.551$ $2.43178$ $0.56$ $2.53449$ $0.565$ $2.95471$
However, the alternative sieve approach may not be the best way to bound $C(\theta)$ when $\theta$ increases. Because of the rapid increase of the size of loss, using the Rosser–Iwaniec sieve is better than using Harman’s sieve when $\theta$ is large. Building on the previous work of Fouvry [12], Baker and Harman [1] showed that
$C(\theta)\leqslant\frac{14}{12-13\theta}-\log\left(\frac{4(1-\theta)}{3\theta}\right)$ (55)
for $\frac{17}{32}\leqslant\theta<\frac{4}{7}$. After calculating, we find that the turning point is between $0.565$ and $0.566$. That is, we use (54) when $\theta<0.565$ and (55) when $\theta\geqslant 0.565$. For $\theta\geqslant\frac{4}{7}$, Baker and Harman [1] also got
$C(\theta)\leqslant\begin{cases}\frac{14}{12-13\theta},&\frac{4}{7}\leqslant\theta<0.6,\\ \frac{8}{3-\theta},&0.6\leqslant\theta\leqslant 0.679.\end{cases}$ (56)
6.11. Proof of Theorem 1
Note that $C(\theta)$ is monotonic increasing, we can use the various upper bounds for $C(\theta)$ in above 10 cases to show that
$\int_{0.5}^{0.679}C(\theta)<0.4993,$ (57)
and Theorem 1 is now proved. The upper constants for the Brun–Titchmarsh theorem with almost all moduli $q\leqslant x^{\theta}$ are given in the above 10 tables.
7. Applications
Clearly our Theorem 1 has many interesting applications (just like the previous Baker–Harman’s result), and we state some of them in this section. Let $g>1$ denote an integer and $f(X)\in\mathbb{Z}[X]$ denote a polynomial of positive degree that has a positive leading coefficient and no multiple roots in its splitting field. Put $u(n)=f\left(g^{n}\right)$. The following two applications focus on the quadratic fields $\mathbb{Q}\left(\sqrt{u(n)}\right)$. Note that this type of quadratic fields was first considered by Shanks [33] with the special case $u(n)=\left(2^{n}+3\right)^{2}-8$. Let $Q_{u}(M,N;s)$ denote the number of integers $n\in[M+1,M+N]$ for which $\mathbb{Q}\left(\sqrt{u(n)}\right)=\mathbb{Q}\left(\sqrt{s}\right)$, where $M,N,s\in\mathbb{Z}$, $N\geqslant 1$ and $s$ square-free. For the case $\text{deg }f\geqslant 3$, Luca and Shparlinski [[26], Theorem 1.3] proved in 2009 that
$Q_{u}(M,N;s)\ll N^{\frac{3}{2(1+\boldsymbol{\Theta})}}(\log N)^{\frac{4+\boldsymbol{\Theta}}{1+\boldsymbol{\Theta}}}.$ (58)
Clearly we have $Q_{u}(M,N;s)\ll N^{\frac{500}{559}+\varepsilon}$ with $\boldsymbol{\Theta}=0.677$ given by Baker and Harman [2] and $Q_{u}(M,N;s)\ll N^{\frac{3}{4}+\varepsilon}$ with conjectured value $\boldsymbol{\Theta}=1$. Note that $\frac{500}{559}\approx 0.895$ and $\frac{3}{4}=0.75$. In 2017, Banks and Shparlinski [[4], Theorem 1.1] improved (58) to
$Q_{u}(M,N;s)\ll N^{\frac{1}{2\boldsymbol{\Theta}}}(\log N)^{2-\frac{1}{\boldsymbol{\Theta}}}.$ (59)
Now we have $Q_{u}(M,N;s)\ll N^{\frac{500}{677}+\varepsilon}$ with $\boldsymbol{\Theta}=0.677$ given by Baker and Harman [2] and $Q_{u}(M,N;s)\ll N^{\frac{1}{2}+\varepsilon}$ with conjectured value $\boldsymbol{\Theta}=1$. Note that $\frac{500}{677}\approx 0.739<0.75$, the unconditional result given by (59) is even better than the conditional result given by (58). As our first application, we apply our Theorem 1 ($\boldsymbol{\Theta}=0.679$) directly to (59) to give that
Theorem 2.
$Q_{u}(M,N;s)\ll N^{\frac{500}{679}}(\log N)^{\frac{358}{679}},\quad\text{where}\quad\frac{500}{679}\approx 0.737.$
We can also obtain the corresponding improved lower bound for $R_{u}(M,N)$, the number of distinct fields among $\mathbb{Q}\left(\sqrt{u(n)}\right)$ as $n$ varies over $M+1,\ldots,M+N$.
In the same paper, Banks and Shparlinski [4] also considered the bound for $Q_{u}(M,N;s)$ on average over square-free $s$. Their [[4], Corollary 1.4] gives
$\sum_{s\leqslant S}Q_{u}(M,N;s)\leqslant\begin{cases}S^{1-\frac{1}{4\boldsymbol{\Theta}}}N^{\frac{1}{2\boldsymbol{\Theta}}+\varepsilon},&S\leqslant N^{\frac{2(1-\boldsymbol{\Theta})}{1+3\boldsymbol{\Theta}}},\\ S^{\frac{1}{2}}N^{\frac{3-\boldsymbol{\Theta}}{1+3\boldsymbol{\Theta}}+\varepsilon},&N^{\frac{2(1-\boldsymbol{\Theta})}{1+3\boldsymbol{\Theta}}}<S\leqslant N^{\frac{4(1-\boldsymbol{\Theta})}{1+3\boldsymbol{\Theta}}},\\ S^{\frac{3}{3+\boldsymbol{\Theta}}}N^{\frac{3-\boldsymbol{\Theta}}{3+\boldsymbol{\Theta}}+\varepsilon},&N^{\frac{4(1-\boldsymbol{\Theta})}{1+3\boldsymbol{\Theta}}}<S\leqslant N^{\frac{2\boldsymbol{\Theta}}{3}},\\ N~(\text{trivial bound}),&N^{\frac{2\boldsymbol{\Theta}}{3}}<S.\end{cases}$ (60)
As our second application, using our Theorem 1 ($\boldsymbol{\Theta}=0.679$) we have
Theorem 3.
$\sum_{s\leqslant S}Q_{u}(M,N;s)\leqslant\begin{cases}S^{\frac{429}{679}}N^{\frac{500}{679}+\varepsilon},&S\leqslant N^{\frac{642}{3037}},\\ S^{\frac{1}{2}}N^{\frac{2321}{3037}+\varepsilon},&N^{\frac{642}{3037}}<S\leqslant N^{\frac{1284}{3037}},\\ S^{\frac{3000}{3679}}N^{\frac{2321}{3679}+\varepsilon},&N^{\frac{1284}{3037}}<S\leqslant N^{\frac{679}{1500}},\\ N,&N^{\frac{679}{1500}}<S.\end{cases}$
Appendix: Shifted primes without large prime factors
Since the shifting process destroys the multiplicative structure of prime numbers, it is conjectured that $P^{+}(p+a)$ has a distribution roughly the same as the distribution of $P^{+}(n)$ over all integers $n\leqslant x$. In fact, one has a special version of this general conjecture:
Conjecture.
There are infinitely many primes $p$ with $P^{+}(p+a)<p^{\boldsymbol{\vartheta}}$ and infinitely many primes $p$ with $P^{+}(p+a)>p^{\boldsymbol{\Theta}}$ for any $0<\boldsymbol{\vartheta},\boldsymbol{\Theta}<1$.
Note that we have already mentioned the historical progress towards the second assertion in the Introduction. For the first assertion, Erdős [10] first proved in 1935 that there exists some $\boldsymbol{\vartheta}<1$ such that $P^{+}(p+a)<p^{\boldsymbol{\vartheta}}$ for infinitely many $p$. After that, many mathematicians improved the upper bound for $\boldsymbol{\vartheta}$. The progression of records for $\boldsymbol{\vartheta}$ can be seen in the following table.
$\boldsymbol{\vartheta}$ $\approx$ Author Year
$2\sqrt{2}-2$ $0.8285$ Wooldridge [35] 1979
$\frac{625}{512e}$ $0.4491$ Pomerance [32] 1980
$\frac{1}{2}\exp\left(-\frac{1}{2.73}\right)$ $0.3465$ Balog [3] 1984
$\frac{1}{2}\exp\left(-\frac{1}{2}\right)$ $0.3033$ Friedlander [13] 1989
$0.2961$ $0.2961$ Baker and Harman [2] 1998
$\frac{15}{32}\exp\left(-\frac{1}{2}\right)$ $0.2844$ Lichtman [23] 2022
However, we cannot improve Lichtman’s result by applying any sieves since we do not have a distribution result for primes in arithmetic progressions with multi-linear forms of moduli larger than $x^{\frac{17}{32}}$.
Acknowledgements
The author would like to thank Igor Shparlinski for providing information about the papers [26] and [4].

### v3

@@
 In 1996, Baker and Harman [1] presented a mixed approach to this problem.
-They combined the traditional average Brun–Titchmarsh approach (with a refined version of [[12], Théorème 3] which will be used in Section 6.10) with the alternative sieve developed by Harman [15], using the estimates of Bombieri–Friedlander–Iwaniec type (see [5] and [6], for example) as arithmetical information inputs.
+They combined the traditional average Brun–Titchmarsh approach (with a refined version of [[12], Théorème 3] which will be used in Section 6.10) with the alternative sieve developed by Harman [15], using the estimates of Bombieri–Friedlander–Iwaniec type (see [5] and [6], for example) as arithmetic information inputs.
 Their method is more powerful when $\theta$ is near to $\frac{1}{2}$.
@@
 Although this result only improved the previous one by $0.001$, it was still a improvement since they used a continuous version of alternative sieve, and any saving on a single point became quite small after integrating.
-Because of the lack of new arithmetical information, even $0.678$ seems impossible without any new things.
-Recently, Maynard [28] got a new breakthrough on the distribution of primes in arithmetical progressions to large moduli $q=q_{1}q_{2}\sim x^{\frac{11}{21}-\varepsilon}$.
-The main theorem in his paper is not applicable here, but we find that one of his lemmas (Proposition 8.3) can be used here as a new arithmetical information input.
+Because of the lack of new arithmetic information, even $0.678$ seems impossible without any new things.
+Recently, Maynard [28] got a new breakthrough on the distribution of primes in arithmetic progressions to large moduli $q=q_{1}q_{2}\sim x^{\frac{11}{21}-\varepsilon}$.
+The main theorem in his paper is not applicable here, but we find that one of his lemmas (Proposition 8.3) can be used here as a new arithmetic information input.
 By inserting this into the sieving process in [1] and [2], together with a careful numerical calculation for sieve integrals, we get the following sharper result.
@@
 The letter $p$, with or without subscript, is reserved for prime numbers.
-We define the boolean function as
-$\texttt{Boole}(\textbf{X})=\begin{cases}1&\text{if }\textbf{X}\text{ is true},\\ 0&\text{if }\textbf{X}\text{ is false}.\end{cases}$ (5)
 2. An outline of the proof
@@
 Note that each $\theta$ corresponds to an independent sieve decomposition.
-In the process of giving asymptotic formulas like (13) in Harman’s sieve, one needs the following types of arithmetical information (with all variables are dyadic constrained, and all coefficients are divisor-bounded):
+In the process of giving asymptotic formulas like (13) in Harman’s sieve, one needs the following types of arithmetic information (with all variables are dyadic constrained, and all coefficients are divisor-bounded):
 $\sum_{q\sim Q}\gamma_{q}\sum_{m}a_{m}\sum_{mn_{1}\cdots n_{j}\in\mathcal{A}^{q}}1=(1+o(1))\frac{1}{\varphi(q)}\sum_{q\sim Q}\gamma_{q}\sum_{m}a_{m}\sum_{mn_{1}\cdots n_{j}\in\mathcal{B}^{q}}1,$ (Type-Ij)
@@
 For the two-dimensional or even three-dimensional alternative sieve, Type-I2 and Type-I3 information is required.
-3. Type-II arithmetical information
+3. Type-II arithmetic information
 In this section we give several results of the form
-$\sum_{q\sim Q}\gamma_{q}\sum_{\begin{subarray}{c}m_{1}\cdots m_{j}\in\mathcal{A}^{q}\\ m_{i}\sim M_{i}\\ 1\leqslant i\leqslant j\end{subarray}}a_{1,m_{1}}a_{2,m_{2}}\cdots a_{j,m_{j}}=(1+o(1))\frac{1}{\varphi(q)}\sum_{q\sim Q}\gamma_{q}\sum_{\begin{subarray}{c}m_{1}\cdots m_{j}\in\mathcal{B}^{q}\\ m_{i}\sim M_{i}\\ 1\leqslant i\leqslant j\end{subarray}}a_{1,m_{1}}a_{2,m_{2}}\cdots a_{j,m_{j}}.$
-The three conditions we want the coefficients to satisfy are as follows:
+$\sum_{q\sim Q}\gamma_{q}\sum_{\begin{subarray}{c}m_{1}\cdots m_{j}\in\mathcal{A}^{q}\\ m_{i}\sim M_{i}\\ 1\leqslant i\leqslant j\end{subarray}}a_{1,m_{1}}a_{2,m_{2}}\cdots a_{j,m_{j}}=(1+o(1))\frac{1}{\varphi(q)}\sum_{q\sim Q}\gamma_{q}\sum_{\begin{subarray}{c}m_{1}\cdots m_{j}\in\mathcal{B}^{q}\\ m_{i}\sim M_{i}\\ 1\leqslant i\leqslant j\end{subarray}}a_{1,m_{1}}a_{2,m_{2}}\cdots a_{j,m_{j}}$
+(The errors, here and in lemmas below, can be bounded by $x(\log x)^{-A}$ for some large $A>1$.) The three conditions we want the coefficients to satisfy are as follows:
 (Condition 3.1: Siegel–Walfisz condition) For any $f\geqslant 1$, $k\geqslant 1$, $b\neq 0$ and $(k,b)=1$, we have
@@
 By (17)–(22), we get the total loss when $\frac{1}{2}\leqslant\theta<\frac{25}{49}$ of
-$\left(\int_{\kappa}^{\frac{\theta}{2}}\int_{1-\theta-t_{1}}^{\theta-t_{1}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{101}]\frac{\omega\left(\frac{t_{2}}{t_{1}}\right)\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}^{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{1}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{2}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{104}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{105}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
-- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in L_{106}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\int_{t_{4}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}}\right$.
-$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{107}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
-$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{108}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+$\left(\int_{(t_{1},t_{2},t_{3})\in L_{101}}\frac{\omega\left(\frac{t_{2}}{t_{1}}\right)\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}^{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\mathbbm{1}_{(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{1}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\mathbbm{1}_{(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{2}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{104}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{105}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
+- $\left(\int_{(t_{1},t_{2},t_{3},t_{4})\in L_{106}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{107}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{108}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
 $=$ $\ E_{101}(\theta)+E_{102}(\theta)+E_{103}(\theta)+E_{104}(\theta)+E_{105}(\theta)-E_{106}(\theta)-E_{107}(\theta)+E_{108}(\theta),$ (22)
@@
 Together, the total loss when $\frac{25}{49}\leqslant\theta<\frac{21}{41}$ can be bounded by
-$\left(\int_{\kappa}^{\frac{\theta}{2}}\int_{1-\theta-t_{1}}^{\theta-t_{1}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{201}]\frac{\omega\left(\frac{t_{2}}{t_{1}}\right)\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}^{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{1}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{2}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{204}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{205}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
-- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in L_{206}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\int_{t_{4}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}}\right$.
-$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{207}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
-$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{208}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{4-7\theta}^{\frac{3}{7}}\texttt{Boole}[(t_{1})\in L_{209}]\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
-$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{3}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{4}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{212}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
+$\left(\int_{(t_{1},t_{2},t_{3})\in L_{201}}\frac{\omega\left(\frac{t_{2}}{t_{1}}\right)\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}^{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\mathbbm{1}_{(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{1}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\mathbbm{1}_{(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{2}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{204}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{205}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
+- $\left(\int_{(t_{1},t_{2},t_{3},t_{4})\in L_{206}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{207}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{208}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1})\in L_{209}}\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
+$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\mathbbm{1}_{(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{3}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\mathbbm{1}_{(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{4}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{212}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
 $=$ $\ E_{201}(\theta)+E_{202}(\theta)+E_{203}(\theta)+E_{204}(\theta)+E_{205}(\theta)-E_{206}(\theta)-E_{207}(\theta)$
@@
 Thus, the total loss when $\frac{21}{41}\leqslant\theta<\frac{16}{31}$ can be bounded by
-$\left(\int_{\kappa}^{\frac{\theta}{2}}\int_{1-\theta-t_{1}}^{\theta-t_{1}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{301}]\frac{\omega\left(\frac{t_{2}}{t_{1}}\right)\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}^{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{1}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{2}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{304}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{305}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
-- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in L_{306}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\int_{t_{4}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}}\right$.
-$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{307}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
-$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{308}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{\tau}^{\frac{3}{7}}\texttt{Boole}[(t_{1})\in L_{309}]\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
-$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{3}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{4}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{312}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
+$\left(\int_{(t_{1},t_{2},t_{3})\in L_{301}}\frac{\omega\left(\frac{t_{2}}{t_{1}}\right)\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}^{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\mathbbm{1}_{(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{1}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\mathbbm{1}_{(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{2}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{304}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{305}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
+- $\left(\int_{(t_{1},t_{2},t_{3},t_{4})\in L_{306}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{307}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{308}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1})\in L_{309}}\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
+$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\mathbbm{1}_{(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{3}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\mathbbm{1}_{(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{4}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{312}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
 $=$ $\ E_{301}(\theta)+E_{302}(\theta)+E_{303}(\theta)+E_{304}(\theta)+E_{305}(\theta)-E_{306}(\theta)-E_{307}(\theta)$
@@
 Combining with the loss integrals from $S_{42}$ and $S_{45}$, the total loss when $\frac{16}{31}\leqslant\theta<\frac{14}{27}$ is no more than
-$\left(\int_{\kappa}^{\frac{\theta}{2}}\int_{1-\theta-t_{1}}^{\theta-t_{1}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{401}]\frac{\omega\left(\frac{t_{2}}{t_{1}}\right)\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}^{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{402}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
-- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in L_{403}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\int_{t_{4}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}}\right$.
-$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{404}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
-$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{405}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{\tau}^{\frac{3}{7}}\texttt{Boole}[(t_{1})\in L_{406}]\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
-$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{3}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{4}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{409}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
+$\left(\int_{(t_{1},t_{2},t_{3})\in L_{401}}\frac{\omega\left(\frac{t_{2}}{t_{1}}\right)\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}^{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{402}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
+- $\left(\int_{(t_{1},t_{2},t_{3},t_{4})\in L_{403}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{404}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{405}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1})\in L_{406}}\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
+$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\mathbbm{1}_{(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{3}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\mathbbm{1}_{(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{4}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{409}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
 $=$ $\ E_{401}(\theta)+E_{402}(\theta)-E_{403}(\theta)-E_{404}(\theta)+E_{405}(\theta)+E_{406}(\theta)+E_{407}(\theta)+E_{408}(\theta)+E_{409}(\theta),$ (33)
@@
 By an exactly same process as the decomposition in Case 4, the total loss when $\frac{14}{27}\leqslant\theta<\frac{11}{21}$ can be bounded by
-$\left(\int_{\kappa}^{\frac{\theta}{2}}\int_{1-\theta-t_{1}}^{\theta-t_{1}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{501}]\frac{\omega\left(\frac{t_{2}}{t_{1}}\right)\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}^{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{502}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
-- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in L_{503}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\int_{t_{4}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}}\right$.
-$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{504}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
-$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{505}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{\tau}^{\frac{3}{7}}\texttt{Boole}[(t_{1})\in L_{506}]\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
-$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{3}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{4}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{509}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
+$\left(\int_{(t_{1},t_{2},t_{3})\in L_{501}}\frac{\omega\left(\frac{t_{2}}{t_{1}}\right)\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}^{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{502}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
+- $\left(\int_{(t_{1},t_{2},t_{3},t_{4})\in L_{503}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{504}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{505}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1})\in L_{506}}\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
+$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\mathbbm{1}_{(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{3}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\mathbbm{1}_{(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{4}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{509}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
 $=$ $\ E_{501}(\theta)+E_{502}(\theta)-E_{503}(\theta)-E_{504}(\theta)+E_{505}(\theta)+E_{506}(\theta)+E_{507}(\theta)+E_{508}(\theta)+E_{509}(\theta),$ (35)
@@
 $\left(\int_{1-\theta}^{\frac{1}{2}}\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
-$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{802}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
-- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in L_{803}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\int_{t_{4}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}}\right$.
-$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{804}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
-$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{805}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\int_{t_{5}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}}\right$.
-$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6})\in L_{806}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{t_{6}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}^{2}}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{802}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
+- $\left(\int_{(t_{1},t_{2},t_{3},t_{4})\in L_{803}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{804}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{805}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6})\in L_{806}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{t_{6}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}^{2}}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
 $+$ $\left(\int_{\tau}^{\frac{3}{7}}\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
@@
 $\left(\int_{1-\theta}^{\frac{1}{2}}\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
-$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{902}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
-- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in L_{903}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\int_{t_{4}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}}\right$.
-$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{904}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
-$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{905}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\int_{t_{5}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}}\right$.
-$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6})\in L_{906}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{t_{6}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}^{2}}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
-$\quad\quad\quad\int_{t_{5}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}}\int_{t_{6}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{2}}$
-$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{907}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
-$\quad\quad\quad\int_{\kappa}^{\min\left(t_{5},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}\right)}\int_{\kappa}^{\min\left(t_{6},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{2}\right)}$
-$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{908}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{909}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
-- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in L_{910}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\int_{t_{4}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}}\right$.
-$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{911}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa^{\prime}}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa^{\prime}}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
-$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{912}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa^{\prime}}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa^{\prime}}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\int_{t_{5}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}}\right$.
-$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6})\in L_{913}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{t_{6}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}^{2}}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa^{\prime}}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa^{\prime}}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
-$\quad\quad\quad\int_{t_{5}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}}\int_{t_{6}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{2}}$
-$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{914}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\frac{1}{2}}\int_{\kappa^{\prime}}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa^{\prime}}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa^{\prime}}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
-$\quad\quad\quad\int_{\kappa^{\prime}}^{\min\left(t_{5},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}\right)}\int_{\kappa^{\prime}}^{\min\left(t_{6},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{2}\right)}$
-$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{915}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{902}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
+- $\left(\int_{(t_{1},t_{2},t_{3},t_{4})\in L_{903}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{904}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{905}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6})\in L_{906}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{t_{6}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}^{2}}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{907}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{908}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{909}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
+- $\left(\int_{(t_{1},t_{2},t_{3},t_{4})\in L_{910}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{911}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{912}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6})\in L_{913}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{t_{6}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}^{2}}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{914}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{915}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
 $+$ $\left(\int_{\tau}^{\frac{3}{7}}\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
@@
 $\left(\int_{1-\theta}^{\frac{1}{2}}\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
-$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{X02}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
-- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in L_{X03}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\int_{t_{4}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}}\right$.
-$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{X04}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
-$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{X05}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\int_{t_{5}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}}\right$.
-$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6})\in L_{X06}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{t_{6}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}^{2}}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
-$\quad\quad\quad\int_{t_{5}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}}\int_{t_{6}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{2}}$
-$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{X07}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
-$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
-$\quad\quad\quad\int_{\kappa}^{\min\left(t_{5},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{2}\right)}\int_{\kappa}^{\min\left(t_{6},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{2}\right)}$
-$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{X08}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{X02}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
+- $\left(\int_{(t_{1},t_{2},t_{3},t_{4})\in L_{X03}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{X04}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{X05}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6})\in L_{X06}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}}{t_{6}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}^{2}}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{X07}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
+$+$ $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5},t_{6},t_{7})\in L_{X08}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}-t_{6}-t_{7}}{t_{7}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}t_{6}t_{7}^{2}}dt_{7}dt_{6}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
 $+$ $\left(\int_{\tau}^{\frac{3}{7}}\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$